### v1

# spacing=nonfrench

Abstract.
Fast and scalable metadata management across multiple metadata servers is crucial for distributed file systems to handle numerous files and directories. Client-side caching of frequently accessed metadata can mitigate server loads, but incurs significant overhead and complexity in maintaining cache consistency when the number of clients increases. We propose FMCache, an in-switch file-system metadata caching framework that leverages programmable switches to serve file-system metadata requests from multiple clients directly in the switch data plane. Unlike prior in-switch key-value caching approaches, FMCache addresses file-system-specific path dependencies under stringent switch resource constraints. We implement FMCache atop Hadoop HDFS and evaluate it on a Tofino-switch testbed using real-world file-system metadata workloads. FMCache achieves up to 181.6% higher throughput than vanilla HDFS and complements client-side caching with additional throughput gains of up to 139.6%. It also incurs low latencies and limited switch resource usage.
1. Introduction
Scaling file-system metadata management across multiple metadata servers is crucial for distributed file systems to handle billions of files and directories. Field studies, both classical (Roselli et al., 2000; Harter et al., 2012) and recent (Wang et al., 2023), indicate that metadata operations dominate file-system requests. For example, 67-96% of file-system requests in Baidu AI Cloud are metadata-related (Wang et al., 2023). As the number of files and directories grows, especially in workloads dominated by small files (Carns et al., 2009; Harter et al., 2014), metadata operations, such as inode lookups, permission checks, and directory traversals, become performance bottlenecks.
Caching frequently accessed metadata on the client side is a common strategy to mitigate loads on metadata servers. File-system metadata access patterns in practice tend to be skewed, with a small fraction of files and directories being accessed far more frequently than others (Leung et al., 2008; Abad et al., 2012; Ananthanarayanan et al., 2011); for example, less than 3% of files account for 34-39% of requests in Yahoo’s HDFS clusters (Abad et al., 2012). However, client-side caching incurs significant overhead in maintaining cache consistency across numerous clients, as metadata updates require notifications to all clients to invalidate cached metadata. Existing client-side caching approaches (Welch et al., 2008; Ren et al., 2014; Li et al., 2017a; Lv et al., 2022) cache limited metadata information to mitigate path resolution overhead (§2.1) but rely on metadata servers for accessing metadata contents, so the performance gains are limited.
Programmable switches (Bosshart et al., 2013) offer a promising alternative by enabling in-switch caching across multiple clients and leveraging a centralized view of requests, so as to eliminate switch-to-server transmissions and server-side processing. While prior work has explored in-switch caching for key-value stores (Li et al., 2016; Jin et al., 2017; Liu et al., 2019; Sheng et al., 2023; Li et al., 2020; Eldakiky et al., 2021; Lee et al., 2021), in-switch caching for file-system metadata poses unique challenges not directly addressed before. First, file-system pathnames, unlike fixed-length keys, have large and variable sizes and cannot readily fit into limited switch resources. Second, accessing a file’s metadata requires accessing its internal directories’ metadata, thereby exacerbating switch resource demands for caching multiple levels of metadata. Third, cache lookups for the metadata of files and directories require multiple iterations under the strict switch programming model, thereby complicating concurrent cache updates and lookups while maintaining consistency and performance.
We propose FMCache, an in-switch file-system metadata caching framework tailored for distributed file systems under skewed, read-intensive workloads. FMCache extends in-switch key-value caches by specifically addressing file-system path dependencies through techniques with performance and correctness guarantees: (i) path-aware cache management to account for path dependencies in cache admission and eviction; (ii) multi-level read-write locking to enable high-performance concurrent cache lookups and updates; and (iii) local hash collision resolution to efficiently and correctly map variable-length paths into fixed-length keys.
We implemented FMCache in P4 (Bosshart et al., 2014) and compiled it into the Tofino switch chipset (tof, [n. d.]; swi, [n. d.]). We integrated FMCache with Hadoop HDFS (hdf, [n. d.]a), while preserving HDFS semantics. We also implemented a state-of-the-art client-side caching approach (Ren et al., 2014; Lv et al., 2022). Our evaluation on four real-world workloads (Ren et al., 2014; Lv et al., 2022; Xu et al., 2024) shows that under 128 simulated metadata servers, FMCache achieves up to 181.6% higher throughput than vanilla HDFS, while client-side caching coupled with FMCache achieves up to 139.6% higher throughput than without FMCache.
2. Background and Motivation
2.1. File-System Metadata Management
File systems organize data in a hierarchical, tree-based namespace, where files and directories are managed as leaf and non-leaf nodes, respectively. Each node is identified by a path and contains metadata (e.g., owner, size, and permission). A path often has multiple internal directories. We refer to a file or directory below the root as the $i$-th level ($i\geq 1$), and the maximum number of levels of a path as the depth. For example, the path /a/b/c.txt has a depth of 3, with levels /a, /a/b, and /a/b/c.txt. In the namespace, we say that path $p$ is an ancestor of path $q$, and $q$ is a descendant of $p$, if $p$ is an internal directory (or prefix) of $q$. We say that $p$ is the parent of $q$, and $q$ is a child of $p$, if $p$ is an ancestor of $q$ with exactly one less level than $q$ (e.g., /a is the parent of /a/b, and /a/b is a child of /a).
In this work, we focus on Hadoop Distributed File System (HDFS) (Shvachko et al., 2010), where a namenode manages the namespace and metadata. HDFS supports various metadata operations for files (e.g., create, delete, open, and close), directories (e.g., mkdir and rmdir), attributes (e.g., chmod, chown, and utime), and data organization (i.e., rename, readdir, and stat). Metadata operations rely on path resolution, which parses and traverses each level of a path to verify metadata for existence and permissions. For scalability, HDFS employs Router-Based Federation (RBF) (hdf, [n. d.]b) to distribute namespace and metadata management across multiple namenodes.
Figure 1. Data plane of a programmable switch.
2.2. Programmable Switches
Programmable switches (Bosshart et al., 2013) enable customized packet processing tasks beyond traditional packet forwarding. We focus on Tofino switches (tof, [n. d.]), which are programmed using the P4 language (Bosshart et al., 2014) and built on the Protocol-Independent Switch Architecture (PISA) based on the Reconfigurable Match-Action Table (RMT) paradigm (Bosshart et al., 2013). We target the generic RMT architecture without relying on any Tofino-specific features, so our design is extensible to other RMT-based switch platforms (bro, [n. d.]; Chole et al., 2017; hua, [n. d.]). A programmable switch comprises a data plane and a control plane. The data plane, as shown in Figure 1, performs packet forwarding via multiple ingress and egress pipelines. Packets enter an ingress pipeline and are forwarded to an egress pipeline via a traffic manager. The control plane manages the data plane by specifying packet processing rules.
Each ingress or egress pipeline processes packets via a series of stages, each of which contains match-action tables (MATs) that execute processing logic. A parser converts packet headers into packet header vectors (PHVs), processed by MATs with on-chip ALUs across stages. Each stage can only access limited SRAM, typically with tens of memory blocks. The switch programming model imposes strict constraints: each memory block can be accessed at most once per PHV traversal, and stages cannot access memory blocks of other stages. After processing, a deparser reconstructs packets. To process packets in multiple iterations, switches support recirculation, which redirects packets from an egress pipeline back to an ingress pipeline. Recirculation should be cautiously used, as it consumes extra switch resources and slows down packet forwarding. In this work, we leverage recirculation for read requests that require in-switch path resolution (§4.1) and write requests that await locks (§5.2).
2.3. Challenges
Designing in-switch metadata caching for distributed file systems poses the following challenges.
Challenge 1: Constrained switch resources. Switches have limited resources and complicate caching implementation. For example, Tofino switches (tof, [n. d.]) provide only 12 stages per ingress or egress pipeline, each processing up to 16 bytes due to ALU word size restrictions. The PHV size is capped (e.g., 768 bytes in Tofino switches (tof, [n. d.])), and all pipelines share limited SRAM. For comparisons, NetCache (Jin et al., 2017) allocates SRAM for access frequency tracking and supports only a maximum value size of 128 bytes for key-value records.
File-system metadata caching is particularly resource-intensive. HDFS file or directory names can reach 255 bytes (def, [n. d.]) and aggravate SRAM and PHV overhead. Path resolution for a path (e.g., /a/b.txt) requires accessing metadata for its ancestors (i.e., / and /a) (§2.1). Caching metadata for all paths and their ancestors is desirable but challenging given limited SRAM. Splitting path processing across multiple levels requires careful synchronization across stages. To mitigate switch resource usage, existing in-switch key-value caches (Jin et al., 2017; Liu et al., 2019; Sheng et al., 2023) offload cache admission and eviction to a centralized controller, but excessive switch-to-controller communications incur high latencies.
Challenge 2: Cache consistency. Enforcing cache consistency under concurrent updates is crucial, as programmable switches handle requests from multiple ingress pipelines and trigger simultaneous cache updates. Cache updates may overlap with path resolution, thereby further complicating cache consistency management. For example, a read request for /a/b.txt requires metadata access for /, /a, and /a/b.txt, while concurrent chmod requests for /a and /a/b.txt update their metadata and may occur between their cache lookups. Thus, the read request can access a mix of pre- and post-updated metadata, leading to inconsistencies.
3. Design Overview
3.1. Goals and Assumptions
FMCache is an in-switch metadata caching framework for distributed file systems, aiming to achieve high throughput and low latencies for file-system metadata requests across multiple metadata servers. It is currently designed for HDFS (§2.1), supporting all HDFS metadata operations via its interface (§7) and preserving HDFS semantics. It is also extensible to other distributed file systems via their respective APIs.
FMCache targets read-intensive file-system metadata workloads, as observed in prior studies (Ren et al., 2014; Lv et al., 2022; Xu et al., 2024). For example, in a LinkedIn HDFS cluster, 84% of 145 million metadata operations are lookups, with only 9% creates and 7% updates (Ren et al., 2014); in Alibaba’s Pangu file system, over 60% of metadata operations are reads (Lv et al., 2022). To ensure consistency, FMCache adopts write-through caching (as in prior in-switch key-value caches (Li et al., 2016; Liu et al., 2017; Jin et al., 2017; Liu et al., 2019)) to update both the in-switch cache and server-side metadata in writes before acknowledging clients. Also, since path resolution always starts at the root, FMCache assumes the root directory’s metadata is persistently cached, with its permissions unchanged and the root not being deleted (Niazi et al., 2017).
3.2. Architecture and Design Roadmap
Figure 2. FMCache’s architecture.
Figure 2 depicts FMCache’s architecture, which comprises both control-plane and data-plane components. In the control plane, a controller manages in-switch cache admission and eviction, while in the data plane, multiple clients send (receive) file-system metadata requests (responses) via the switch to (from) multiple metadata servers (or servers for short). FMCache employs three techniques to enable efficient in-switch file-system metadata caching.
Path-aware cache management (§4). FMCache caches file-system metadata in the switch data plane. FMCache hashes variable-length paths into fixed-length keys and manages file-system metadata in key-value pairs as in existing in-switch key-value caches (Li et al., 2016; Jin et al., 2017; Liu et al., 2019; Sheng et al., 2023). But unlike in-switch key-value caches, FMCache addresses path dependencies across hierarchical levels during cache admission and eviction.
Multi-level read-write locking (§5). To synchronize concurrent metadata operations, FMCache adopts multi-level read-write locking using (i) lock counters as read-write locks for different path levels and (ii) validation flags for verifying the validity of cached paths. It ensures reliable locking and unlocking during path resolution even in unreliable networks.
Local hash collision resolution (§6). Mapping paths to fixed-size keys in the switch data plane can cause hash collisions, leading to incorrect metadata retrieval. For example, if two paths /a and /b share the same hash and /a is cached, a request for /b can erroneously return /a’s metadata. FMCache proposes a local hash collision resolution approach, where the controller assigns unique values, called tokens, to hash-colliding paths and distributes the tokens to clients, servers, and the switch, so as to allow local resolution of hash collisions without the controller’s intervention for every request. FMCache ensures no incorrect metadata retrieval.
4. Path-Aware Cache Management
FMCache incorporates path awareness into cache admission and eviction by caching frequently accessed paths and their ancestors, so as to mitigate cache misses during path resolution. This ensures that if a path is cached, its ancestors are also cached, thereby allowing the reuse of cached metadata across different paths with the common ancestors.
4.1. Path Representation
FMCache maps a file-system pathname into fixed-size hash keys for efficient switch operations. For a read operation, a FMCache client partitions a path into multiple levels (including the root) and computes a hash for each level. For example, /a/b/c.txt is partitioned into /, /a, /a/b, and /a/b/c.txt, with each level being hashed. To avoid redundant computations, the hash of the root directory / is pre-computed and cached on the client side. For a write operation, the client computes a hash for the complete path without partitioning, as all writes are forwarded to the server (that holds the namespace) for path resolution under write-through caching (§3.1). Currently, FMCache uses the first 64 bits of a 128-bit MD5 hash as the hash value.
FMCache caches file-system metadata as key-value records in the switch. Each record is identified by its hash key, and its value contains file or directory metadata. Files and directories have common metadata fields (i.e., type, permissions, owner/group IDs, and timestamps); while files additionally include the size and replication factor fields. Each file’s metadata has 40 bytes, and each directory’s has 24 bytes.
FMCache manages cache across multiple stages in the switch. It leverages MATs across two stages in a single ingress pipeline to perform cache lookups based on hash keys, so as to avoid cross-pipeline synchronization (§5) and enforce logical dependency (i.e., cache hit/miss status is resolved before metadata retrieval can proceed). Since packets are strictly processed in ingress-to-egress order, this placement enables fast cache status checks. The switch manages metadata in 32 registers across eight stages in the egress pipelines (§7). To simplify routing, the switch caches metadata in the egress pipeline that is connected to the corresponding server.
Clients issue metadata requests, including hash keys and full pathnames, to the switch, which reports frequently accessed pathnames to the controller during cache admission (§4.2). For reads, the switch performs path resolution by issuing cache lookups and permission checks for each level, starting from the root, and recirculates the request to the ingress pipeline for each next level. For a cache hit (i.e., the metadata for the path and all its ancestors are cached), the switch returns the metadata if all levels’ permission checks pass, or an error if any check fails. For a cache miss, the switch forwards the request to the server. Due to limited stages and memory access constraints (§2.2), FMCache cannot perform path resolution for multiple path levels in a single pass, but instead leverages recirculation for path resolution while limiting overhead (§8.2).
4.2. Cache Admission and Eviction Workflows
Data structures. FMCache monitors path access frequencies for read operations (excluding access to ancestors during path resolution). For uncached paths, which dominate path access traffic, FMCache uses a Count-Min Sketch (CMS) (Cormode and Muthukrishnan, 2005), as in prior studies (Jin et al., 2017; Liu et al., 2019; Sheng et al., 2023), to estimate their access frequencies within fixed-size memory with provable error bounds. For cached paths, FMCache tracks their exact access frequencies using a frequency counter array. The switch periodically reports the access frequencies of all cached paths to the controller for eviction decisions, and resets the CMS and frequency counter array after each reporting.
Cache admission. FMCache’s cache admission is triggered by the switch data plane. During runtime, the switch monitors the access frequencies of uncached paths via the CMS and identifies hot paths (i.e., those exceeding a pre-defined CMS threshold) for admission. When a hot path $p$ is detected, the switch notifies the controller, which retrieves the metadata for $p$ and its uncached ancestors from the servers. The controller communicates with servers using UDP (and issues retransmissions if needed), bypassing the switch data plane to avoid critical-path overhead, as in prior work (Jin et al., 2017; Sheng et al., 2023).
The controller verifies cache capacity. If the cache is not full, the controller sends the hash keys and metadata for $p$ and its uncached ancestors to the switch for admission (note that the controller maintains a global view of all cached paths). Unlike NetCache (Jin et al., 2017) (without path awareness), FMCache admits the metadata for both $p$ and its uncached ancestors to ensure path-aware caching.
Cache eviction. When the cache is full and an uncached hot path $p$ is reported, the controller triggers cache eviction to reclaim cache space for $p$ and its uncached ancestors. It selects cached path candidates for eviction based on periodically reported access frequencies, prioritizing the least frequently accessed path with no cached descendants. If the selected path is the only cached child of its parent, the parent is also selected. The controller recursively includes ancestors as candidates until reaching an ancestor with multiple cached children or the root. The rationale of this recursive inclusion is that the selected ancestors are unlikely accessed by themselves, as directory operations are infrequent in practice (Ren et al., 2014; Lv et al., 2022; Xu et al., 2024) (e.g., only 4.2% of metadata operations in Alibaba’s Pangu file system are directory-related (Lv et al., 2022)).
The controller selects candidates up to a pre-defined threshold, currently set as twice the number of paths to be admitted (i.e., $p$ and its uncached ancestors), as the periodically reported access frequencies may vary over time and selecting more candidates than being admitted can avoid mistakenly evicting hot paths. It reloads the current access frequencies of the selected candidates from the switch and evicts the least frequently accessed path with no cached descendants, along with any ancestors having only one cached child. It repeats the selection until sufficient cache space is reclaimed. Finally, the controller notifies the switch to evict the selected paths and admits $p$ and its uncached ancestors.
Figure 3. Example of cache admission and eviction workflows.
Example. Figure 3 depicts the cache admission and eviction workflows. Consider a full cache with five records, holding /, /a, /e, /a/b.txt, and /e/f.txt with access frequencies 0, 1, 1, 12, and 5, respectively, where / is always cached. Suppose that the switch reports an uncached hot path /c/d.txt, with an access frequency 10, to the controller. As /c is uncached, the controller aims to admit both /c and /c/d.txt. With a full cache, the controller triggers cache eviction to select four candidates (i.e., /a, /e, /a/b.txt, and /e/f.txt), twice the number of paths to be admitted, and reloads their current access frequencies (e.g., 0, 0, 5, and 10, respectively). The path /a/b.txt has the lowest access frequency among all paths without cached descendants, and its ancestor /a (with only one cached child) will also be evicted. The controller notifies the switch to evict /a and /a/b.txt, and admits /c and /c/d.txt.
5. In-Switch Read-Write Locking
FMCache adopts multi-level read-write locking to ensure cache consistency under concurrent metadata updates. The locking mechanism operates entirely within the switch data plane to avoid controller overhead.
5.1. Lock Design
FMCache employs multiple lock counter arrays and a validation array, both implemented as register arrays in the switch data plane and initialized with zero entries. Each cached path is associated with one lock counter (a slot in a lock counter array) and one validation flag (a slot in the validation array), indexed by the path’s hash key.
Lock counter arrays. Each lock counter array corresponds to specific path levels. Each of its lock counters records the number of active read requests for a cached path at that level. Due to switch resource constraints, FMCache allocates eight counter arrays of 65,536 16-bit counters each, while each counter supports up to 65,535 concurrent read requests. The first seven arrays are assigned to levels 1 through 7 (e.g., /a for level 1 (§2.1)), while the eighth array handles all remaining deeper levels (using the hash key of level 8). This design maximizes concurrent access to shallow paths, based on empirical evidence that most metadata requests are aggregated at small depths (Agrawal et al., 2007; Meyer and Bolosky, 2012; Douceur and Bolosky, 1999); for example, 90% (Agrawal et al., 2007; Meyer and Bolosky, 2012), or even close to 100% (Douceur and Bolosky, 1999), of accessed paths have a depth of no more than 10. FMCache maps a cached path to a lock counter array based on its level, and uses its hash key’s last 16 bits to associate the path with a specific lock counter.
Validation array. The validation array (Jin et al., 2017; Sheng et al., 2023) tracks metadata validity for all cached paths. A validation flag of one indicates valid metadata and allows reads from the cache, while a flag of zero indicates invalid or updating metadata and all reads are directed to servers. After a cache update, its validation flag is set to one to permit subsequent reads.
5.2. Read and Write Flows
FMCache places the validation array across all egress pipelines, co-located with value register arrays (§4.1) for efficient validity checks. Lock counter arrays reside in a single ingress pipeline to avoid cross-pipeline synchronization. FMCache redirects requests arriving at other ingress pipelines to the ingress pipeline holding lock counter arrays via cross-pipeline recirculation (§7). This placement is critical, as lock counter arrays should be positioned before the validation array to ensure correctness, and they cannot be placed in egress pipelines due to insufficient switch resources. In FMCache’s deployment, the switch data plane is not a bottleneck, and consistent packet processing ordering is maintained across ingress and egress pipelines, so FMCache ensures efficient and correct locking and validation operations.
Reads. Reads are classified as (i) single-path reads (e.g., stat), which retrieve only metadata of the requested path; and (ii) multi-path reads (e.g., readdir), which retrieve metadata of the requested path and its descendants. Single-path reads are served from the in-switch cache, while multi-path reads are forwarded to servers to ensure correctness, as descendant paths may be uncached and servers maintain the authoritative namespace to resolve partially caching scenarios. Multi-path reads are rare in practice (e.g., only 3.9% in Alibaba’s workloads (Lv et al., 2022); see Table 1). Upon receiving a (single-path) read request, the switch checks if the path’s last level is cached, implying that all its ancestors are also cached (§4). If so, the switch increments the lock counter for each path level by one and resolves the path via recirculation.
During path resolution, the switch checks the validation flag for each level. If the validation flag is one (i.e., valid metadata and no ongoing write), the switch retrieves the metadata from cache and performs permission checks. If the permission checks pass, the switch proceeds to the next level and decrements the lock counter for the previous level by one. After resolving the whole path, the switch decrements the lock counter for the last level by one, ensuring that all lock counters are released. If permission checks fail at any level, the switch sends an error response to the client and decrements the lock counters from the failure point to the requested path by one. Conversely, if the validation flag is zero (i.e., invalid or updating metadata), the switch forwards the read request to the server, which returns a response. The switch then decrements all lock counters from the invalid metadata point to the requested path by one, and returns an ACK to the server. If the server does not receive the ACK before timeout, it retransmits the same response.
Any ACK loss can cause duplicate lock decrements via retransmissions and violate correctness. FMCache handles ACK loss via a server-switch sequence-number protocol. Each server tags a lock-related response with a local sequence number (initially 0), and increments the sequence number upon receiving an ACK from the switch. The switch also maintains per-server expected sequence numbers in a sequence counter array. Upon receiving a server’s response, the switch compares the embedded sequence number with its expected value: if they match, the switch increments the expected value (for the next response), processes lock updates, forwards the response to the client, and returns an ACK to the server; if the embedded sequence number is lower, it indicates a duplicate, so the switch only sends an ACK to the server to suppress further retransmissions.
Figure 4. Example of processing a read request under multi-level read-write locking.
For example, consider three cached paths /a, /a/b, and /a/b/c.txt. For a read request open /a/b/c.txt, the switch maps each path level /a, /a/b, and /a/b/c.txt to a lock counter in the first, second, and third lock counter arrays, respectively, with corresponding validation flags (see Figure 4). The switch increments the lock counters for all levels by one. If /a’s validation flag is one and its permission check passes, it processes /a/b and decrements /a’s lock counter by one. If /a/b’s validation flag is zero, it forwards the request to the server, which retrieves /a/b/c.txt’s metadata. The server returns a response, which triggers the switch to decrement the lock counters for both /a/b and /a/b/c.txt by one (assuming no packet loss). The switch then returns an ACK to the server.
Writes. Writes are classified as (i) single-path writes (e.g., chmod and chown), which update only the requested path, and (ii) multi-path writes (e.g., chmod -r and chown -r), which update the requested path and its descendants. The locking mechanism ensures consistency for both.
(i) Single-path writes. For a single-path write, the switch checks if the path is cached. If so, it checks the corresponding lock counter based on its level and hash key. If the lock counter is non-zero (i.e., with ongoing reads), the switch recirculates the request until the lock counter reaches zero (i.e., no ongoing read). Then, the switch sets the validation flag to zero and forwards the request to the server. If the write is successfully completed, the server returns a response, which triggers the switch to update the cached path’s metadata and set the validation flag to one; otherwise, only the validation flag is set to one without cache updates.
Note that a write delayed by ongoing reads may be overtaken by a new read arriving before recirculation, so it is theoretically possible for a write to be blocked indefinitely by continuous reads. However, the short recirculation time makes this unlikely. FMCache’s locking ensures correctness and allows the write to proceed after sufficient recirculations.
(ii) Multi-path writes. Multi-path writes follow the same single-path write processing until the request reaches the server. If the write is successfully completed, the server updates the cache for all cached descendants before the requested path, so that the requested path remains invalidated until all cached descendants are fully updated. By performing path resolution in a top-down manner, FMCache prevents reads from accessing cached descendants before the requested path is updated (i.e., until the cache updates for a multi-path write are completed), so as to ensure cache consistency.
6. Local Hash Collision Resolution
Figure 5. Example of token generation and distribution. Note that the controller also assigns tokens for $p$’s uncached ancestors in Steps 2-4. We omit the details for brevity.
FMCache maps file-system pathnames to 64-bit hash keys (§4.1), so hash collisions are possible (albeit unlikely) and lead to incorrect metadata retrieval. While the controller holds a global view of all cached paths (§4.2), querying the controller to resolve hash collisions for every request is impractical due to high switch-to-controller latencies (§2.3). FMCache adopts a local, token-based hash collision resolution mechanism by synchronizing the controller’s global view of cached paths with the switch, clients, and servers, so as to ensure correctness without compromising performance.
Hash collision resolution with tokens. A token is an 8-bit value paired with a 64-bit hash key to uniquely identify a path. Valid tokens range from 1 to 255, while 0 indicates invalid. A cached path is assigned token 1 if no collision occurs, or the next available token (e.g., 2) if it collides with an existing cached path. Thus, we can associate valid tokens with up to 255 paths with the same colliding hash key. The probability of two paths colliding is $1/{2}^{32}$ based on the birthday paradox (Flajolet et al., 1992), making it unlikely for more than 255 paths to share the same hash key in a large namespace. Thus, 8-bit tokens suffice for hash collision resolution.
FMCache maintains two unordered map structures: (i) a path-token map, which records paths and their tokens (path-token pairs), and (ii) a hash-token map, which records hash keys and their tokens (hash-token pairs). The controller keeps both maps, each client and server holds a path-token map, and the switch holds a hash-token map.
Token generation and distribution. During cache admission, the controller assigns tokens to each level of a hot path, as shown in Figure 5. When the switch reports a hot path $p$ for admission ( 1), the controller checks its map structures. If $p$ is the first time being admitted and its hash key is unique, the controller assigns a token of value one; if a collision occurs, it assigns the next available token (e.g., $t$) ( 2). The controller updates its path-token and hash-token maps with the new path-token and hash-token pairs, respectively ( 3). These entries persist in the controller’s maps even after cache eviction to ensure consistency. If an evicted path is later re-admitted, its previously assigned token is reused. In the worst case, the controller may store entries for all paths ever cached. For scalability, the controller can use persistent key-value stores (e.g., RocksDB (roc, [n. d.])) for the map structures.
The controller distributes each admitted path and its token to the switch and the relevant server that holds the path ( 4). The switch adds the hash key and token to the hash-token map, while the server adds the full path and token to the path-token map. These entries are removed during cache eviction, as notified by the controller.
Figure 6. Example of how a client updates its path-token map. Note that the client also attaches tokens of $p$’s ancestors in Step 1, the server returns the tokens for $p$’s ancestors in Step 5, and the client adds the tokens in Step 6. We omit the details for brevity.
Token attachment to requests. Clients update their path-token maps when issuing read or write requests, as shown in Figure 6. Initially, with an empty path-token map, a client attaches an invalid token (value zero) to a request (e.g., reading path $p$’s metadata) sent to the switch ( 1). The switch detects a cache miss by querying its hash-token map ( 2) and forwards the request to the server ( 3). The server retrieves the metadata, checks its path-token map ( 4), and responds with the metadata and valid tokens ( 5). The client updates its path-token map with the valid tokens for future requests ( 6). For cached paths, the switch confirms cache hits when valid tokens are included in the request. To manage token accumulation, clients can attach an expiry time (e.g., one hour) to path-token map entries and clear expired entries.
Discussion. FMCache’s token-based mechanism ensures correctness with modest client complexity and overhead. Clients must maintain path-token maps and attach tokens to requests, and incur a cache miss on the first read of a cached path due to the initial lack of a valid token. Also, including tokens in requests adds minor communication overhead. Theoretically, the number of tokens grows linearly with the depth of a path. Nevertheless, as most paths have small depths (§5.1), the communication overhead is minimal.
7. Implementation
We have built a prototype of FMCache, including the controller, in-switch cache, clients, and servers.
Controller. The controller is implemented in C++ with 2.8 K LoC. It leverages APIs provided by the Tofino switch compiler (tof, [n. d.]) to interact with the switch data plane, including configurations and updates of MATs and registers, for cache admission and eviction (§4) and token management (§6).
In-switch cache. The in-switch cache is implemented in P4 with 5.1 K LoC and compiled for a Tofino switch (tof, [n. d.]; swi, [n. d.]). It comprises several components: (i) a hash-token map using MATs for cache lookups and local hash collision resolution, where each entry is a 9-byte key (8-byte hash key and 1-byte token); (ii) 32 register arrays of 32-bit slots for storing cache entries; (iii) a three-row CMS with three register arrays, each with 64K 16-bit slots; (iv) a frequency counter array with a register array of 32-bit slots; (v) lock counter arrays with eight register arrays, each with 16K 16-bit slots; (vi) a validation array with a register array of 1-bit slots; and (vii) a sequence counter array with a register array of 8-bit slots.
Client implementation. The C++ client driver supports multi-threaded workload execution, where each thread simulates a logical client. Each client maintains a path-token map for local hash collision resolution (§6). It communicates with the switch via UDP-encapsulated packets customized for metadata requests, with re-transmission support for reliability. We integrate the driver with the mdtest (mdt, [n. d.]) benchmarking tool for HDFS metadata operations.
Server implementation. Each server is implemented in C++ with 3 K LoC and hosts an HDFS (v3.2) namenode (hdf, [n. d.]a). We use RBF (hdf, [n. d.]b) with the HASH_ALL policy, which uses consistent hashing to distribute files evenly across all namenodes for load balancing and creates directories on all namenodes. Each server manages a subset of the metadata namespace, connects to its local namenode via the C++ HDFS client library libhdfs3 (lib, [n. d.]), and serves client requests while maintaining a path-token map for local hash resolution (§6).
8. Evaluation
8.1. Methodology
Workload Operation ratio Read ratio
Alibaba (Lv et al., 2022) 52.6% open/close, 9.59% create, 3.9% readdir, 69.1%
0.1% chmod, 11.9% delete, 12.4% stat, 0.2% statdir,
0.005% mkdir, 0.005% rmdir, 9.3% file rename
Training (Xu et al., 2024) 54.32% open/close, 28.5% stat, 0.13% readdir, 83.1%
9.01% create, 0.13% mkdir, 0.13% rmdir,
9.01% delete, 0.13% statdir
Thumb (Xu et al., 2024) 57.01% open/close, 28.44% stat, 0.13% readdir, 85.7%
14.16% create, 0.13% mkdir, 0.13% statdir
LinkedIn (Ren et al., 2014) 84% open/getattr, 9% create/mkdir, 84%
7% chmod/delete/rename
Table 1. Real-world workloads and their metadata operation ratios.
Testbed. We evaluate FMCache on a testbed comprising a 3.2 Tbps two-pipeline Tofino switch (tof, [n. d.]; swi, [n. d.]) and three physical machines. Two machines host servers, each with a 2.40 GHz 10-core Intel Xeon Silver 4210R CPU, 128 GiB DRAM, and a 2 TB HDD (Dell PERC H330 Mini). The client driver runs on the remaining machine, with a 2.40 GHz, 16-core Intel Xeon Silver 4314 CPU, 128 GiB DRAM, and a 960 GB NVMe SSD (Micron 9300 PRO). Each machine is connected to the switch via a 40 Gbps NIC (Mellanox ConnectX-5 CX516A). The client machine uses one pipeline, and the two server machines use another pipeline. All counter arrays (§7) reside in the server-connected pipeline.
Since our two-pipeline Tofino switch lacks native support for cross-pipeline recirculation, we physically connect the designated ingress pipeline hosting the lock counter arrays to another ingress pipeline using a physical wire (Sheng et al., 2023). This enables requests from another ingress pipeline to be recirculated to the designated ingress pipeline (§5). Future programmable switches with native cross-pipeline recirculation would eliminate this requirement.
Workloads. We evaluate FMCache using four real-world workload traces: (i) Alibaba’s Pangu file-system instances (Alibaba) (Lv et al., 2022), (ii) convolutional neural network training (Training) (Xu et al., 2024), (iii) the processing of one million thumbnail images (Thumb) (Xu et al., 2024), and (iv) a LinkedIn HDFS cluster (LinkedIn) (Ren et al., 2014). Table 1 summarizes the proportion of metadata operations and the read ratio for each workload.
We refine the workloads for our evaluation. To focus on metadata performance, we exclude reads and writes of file data, following prior studies (Niazi et al., 2017; Li et al., 2017a; Ren et al., 2014; Lv et al., 2022). For Training and Thumb that include file reads and writes (Xu et al., 2024), we exclude these operations and normalize the remaining metadata operations. Since we exclude file writes, we treat close as a read operation, while it updates both modification and access timestamps if the closed file has been updated. HDFS updates access timestamps hourly by default (def, [n. d.]), and we exclude timestamp updates as they are infrequent.
For LinkedIn, as the original paper (Ren et al., 2014) does not provide operation ratios, we adjust the ratios based on trace analysis (Lv et al., 2022) as follows: open (42%), getattr (42%), create (4.5%), mkdir (4.5%), chmod (1%), delete (3%), and rename (3%). We assign the smallest ratio to chmod as it is less frequent than delete and rename (Lv et al., 2022). Further, getattr corresponds to stat and statdir in HDFS. We replace getattr with stat, as file operations are much more frequent than directory operations (Lv et al., 2022), and focus on file renaming for rename.
Generation of metadata operations. We use mdtest (mdt, [n. d.]) to generate file-system namespaces and metadata operations. By default, we configure a path depth of nine, as metadata requests are often aggregated at small depths (Agrawal et al., 2007; Meyer and Bolosky, 2012; Douceur and Bolosky, 1999)5.1). We create 32 million empty files to focus on metadata performance; empty files are also used in prior evaluation (Ren et al., 2014; Xiao et al., 2015; Niazi et al., 2017; Li et al., 2017a; Lv et al., 2022; Wang et al., 2023). To simulate workload skewness, we use the 80/20 rule, with 80% of operations on 20% of directories and files (Wang et al., 2021a; Xu et al., 2024). We model file access frequencies across levels using a power-law distribution with an exponent of 0.9 and randomly assigning an operation type using the ratios in Table 1. For statdir and readdir, we choose the parent directory of the selected file. For mkdir and rmdir, we use separate directories to avoid removing non-empty directories. We address various workload settings in §8.3.
We mix all metadata operations in each workload. However, rename, delete, and rmdir inherently involve metadata modifications that necessitate the granting and revoking of leases in HDFS (Wang et al., 2009). The lease-based operations on frequently created and deleted files can slow down all metadata operations when we perform stress tests. Thus, we place rename, delete, and rmdir at the end of the request sequence, so as to keep metadata operations at high rate.
Scaling servers. Although our testbed contains only two physical servers, our Tofino switch has significantly higher forwarding throughput (3.2 Tbps (swi, [n. d.])) than server throughput (tens of KOPS) and is not our evaluation bottleneck. To simulate larger-scale deployment with limited hardware while maintaining realism in evaluation, we adopt the well-known server rotation approach as in prior in-switch caching studies (Jin et al., 2017; Sheng et al., 2023). Let $N$ be the number of simulated servers. We assign files to $N$ simulated servers using the HDFS’s RBF policy (§7). We calculate each server’s load from file access frequencies, and identify the bottlenecked server with the highest load. We then perform our evaluation in $N$ iterations. We first deploy the bottlenecked server in one physical machine and fully saturate it to measure its performance. In the following N1N-1 iterations, we reset the physical machines to the initial state, pair the bottlenecked server with one of the N1N-1 non-bottlenecked servers, deploy them on the two physical machines, and measure the performance of the non-bottlenecked server. We aggregate the performance of all servers as the overall performance. By default, we simulate 16 servers and increase the scale to 128 servers (Exp#1).
Baselines. We consider client-side caching (CCache) by faithfully following the state-of-the-art client-side caching implementations in IndexFS (Ren et al., 2014) and InfiniFS (Lv et al., 2022)9). CCache uses RocksDB (v6.22.1) as a key-value store to keep all metadata instead of in an HDFS namenode in each simulated server to eliminate HDFS’s path resolution overhead. Each CCache client locally caches directories’ permission metadata, and forwards all read requests to servers for attribute retrieval. We implement lazy invalidation (Lv et al., 2022) for cache consistency, which outperforms lease-based cache management (Ren et al., 2014). We do not compare FMCache with in-switch key-value caches (e.g., NetCache (Jin et al., 2017) and FarReach (Sheng et al., 2023)) since they are designed for key-value stores and do not support file-system semantics (i.e., they cannot ensure correctness and consistency for file-system operations).
We evaluate four schemes: (i) NoCache, which does not employ any caching and performs metadata operations directly with HDFS namenodes in all simulated servers; (ii) CCache, our client-side caching implementation; (iii) FMCache, which extends NoCache with in-switch caching; and (iv) FMCache+, which extends CCache with in-switch caching.
Before each experiment, we pre-load 32 million files into HDFS namenodes for NoCache and FMCache, and the corresponding paths and metadata into RocksDB for CCache and FMCache+. We pre-load the 5,000 hottest files and their ancestors into the in-switch cache for FMCache and FMCache+. We set the pre-defined CMS threshold of FMCache and FMCache+ to 10 and 20, respectively (a larger threshold is used for FMCache+ due to its higher throughput with client-side caching), and reset the CMS and the frequency counter array every two seconds. We simulate 128 client threads, which sufficiently saturate back-end servers. The client-side cache of each simulated client in CCache and FMCache+ is allocated 4 MiB (Lv et al., 2022). We plot the average results over five runs, with error bars as 95% confidence intervals under the Student’s t-distribution.
8.2. Performance Analysis
Refer to caption Refer to caption
(a) 16 servers
(b) 128 servers
Figure 7. (Exp#1) Performance under real-world workloads.
(Exp#1) Throughput under real-world workloads. Figure 7 shows the throughput results on 16 and 128 simulated servers via server rotation under four real-world workloads. Under 16 servers, FMCache increases NoCache’s throughput by 22.9%, 51.7%, 53.7%, and 47.6% in Alibaba, Training, Thumb, and LinkedIn, respectively, due to load balancing via in-switch caching, while NoCache suffers from load imbalance.
FMCache has lower throughput than CCache by around 38.1% for all workloads. The reason is that CCache reduces the path resolution overhead of HDFS by caching directories’ permission metadata in clients and bypasses the overhead of lease management and distributed transactions in HDFS. However, by integrating FMCache with CCache, FMCache+ increases CCache’s throughput by 8.0%, 33.5%, 36.3%, and 31.5% in Alibaba, Training, Thumb, and Linkedin, respectively, as CCache forwards all read requests to servers for attribute retrieval (§8.1) while FMCache+ improves load balancing via in-switch caching. Our results show that FMCache complements client-side caching to improve metadata performance.
Under 128 servers, FMCache increases NoCache’s throughput by 11.0%, 134.6%, 181.6%, and 71.2% in Alibaba, Training, Thumb, and LinkedIn, respectively. FMCache+ increases CCache’s throughput by 14.7%, 103.6%, 139.6%, and 57.3% for the same workloads, respectively. FMCache and FMCache+ achieve significantly higher throughput gains than that in 16 servers (except in Alibaba) as they improve server scalability via load balancing. The throughput gains from 16 to 128 servers are marginal in Alibaba, which has the largest write ratio among the four workloads. Since FMCache and FMCache+ adopt write-through caching, maintaining cache consistency for extensive writes incurs substantial overhead.
(Exp#2) Single-operation performance. Figure 8 shows the throughput of individual metadata operations: open, stat, create, mkdir, rename, chmod, delete, and rmdir. For read operations (open and stat), FMCache increases NoCache’s throughput by 80.1% and 80.5%, while FMCache+ increases CCache’s throughput by 74.9% and 77.5%, respectively. For write operations (create, mkdir, rename, chmod, delete, and rmdir), FMCache has lower throughput than NoCache by 2.7%, 0.2%, 13.2%, 36.5%, 12.7%, and 14.6%, respectively, while FMCache+ has lower throughput than CCache, by 4.9%, 3.5%, 7.2%, 12.3%, 8.1%, and 6.4%, respectively. The performance drops stem from the switch’s cache maintenance overhead. Among write operations, chmod incurs the highest overhead as chmod on cached paths requires fetching metadata from HDFS namenodes and updating the in-switch cache with the latest metadata. In contrast, rename, delete, and rmdir on cached paths only mark the paths as deleted in the in-switch cache, while create and mkdir only involve creations on new files and directories on HDFS namenodes, respectively, and will not trigger cache updates.
Figure 8. (Exp#2) Single-operation performance.
(Exp#3) Impact of chmod ratio. We further analyze chmod, which triggers frequent cache updates and shows the most performance drops in FMCache and FMCache+ (Exp#2). We generate mixed read-write workloads composed of open (i.e., reads) and chmod (i.e., writes) with different ratios. Each of open and chmod follows a power-law distribution with an exponent of 0.9. Figure 9 shows that at 0% chmod ratio, FMCache and FMCache+ achieve higher throughput than NoCache and CCache, respectively, but as the chmod ratio increases, their throughput decreases, while the throughput of NoCache and CCache remains stable. FMCache and FMCache+ begin to show throughput degradations when the chmod ratio exceeds 50%. At 100% chmod ratio, FMCache and FMCache+ reach the lowest throughput, 36.5% and 12.3% lower than NoCache and CCache, respectively. Nevertheless, real-world workloads have low chmod ratios (Table 1), so FMCache and FMCache+ still maintain performance gains in practice (Exp#1).
FMCache’s multi-level read-write locking design (§5) significantly reduces the overhead in writes. We compare multi-level locking (MultiLock) with single-level locking (SingleLock), which always maps a full path to the first lock counter array. We measure the recirculation count (i.e., number of recirculations per request) by the monitoring tool Barefoot Shell (bfs, [n. d.]). We do not compare MultiLock and locking-disabled in-switch caches, which cannot maintain cache consistency. Table 2 shows that for read-only and write-only workloads, the recirculation counts for both SingleLock and MultiLock are 7.63 and 1, respectively, as FMCache recirculates a read request multiple times for path resolution and a write request once for lock access. MultiLock significantly reduces the recirculation count of SingleLock (e.g., by 93.1% for 75% chmod) due to mitigated lock contention.
Figure 9. (Exp#3) Throughput versus chmod ratio.
chmod ratio SingleLock MultiLock
0% 7.63 7.63
25% 442.06 37.21
50% 436.29 39.42
75% 369.43 25.67
100% 1 1
Table 2. (Exp#3) Recirculation count versus chmod ratio.
(Exp#4) Latency analysis. We analyze request latencies by adjusting the request sending rate to a target throughput, as in prior studies (Cheng et al., 2015; Jin et al., 2017; Sheng et al., 2023; Didona and Zwaenepoel, 2019), so as to analyze the trade-off between latencies and throughput. We focus on (i) a read-only workload that issues 32 million open requests and (ii) the Alibaba workload, which has the largest write ratio (Table 1). The two workloads show FMCache’s best- and worst-case performance, respectively, under write-through caching (§3.1). We follow a power-law distribution with an exponent of 0.9 and consider 16 simulated servers.
Refer to caption Refer to caption
(a) Read-only workload
(b) Read-only workload
Refer to caption Refer to caption
(c) Read-only workload
(d) Alibaba
Refer to caption Refer to caption
(e) Alibaba
(f) Alibaba
Figure 10. (Exp#4) Latency analysis.
Figure 10 shows the average, p95, and p99 latency results. For the read-only workload, all schemes show low latencies at low target throughput (less than 0.1 MOPS) as the servers are not saturated and do not have queueing delays. When the target throughput increases to 0.2 MOPS, NoCache starts to show increasing latencies due to load imbalance under higher loads. FMCache maintains low latencies via in-switch caching. For example, at target throughout 0.26 MOPS, FMCache reduces NoCache’s average, p95, and p99 latencies by 64.6%, 89.8%, and 87.7%, respectively. The same trend is also observed for FMCache+ and CCache. For example, at target throughput 0.59 MOPS, FMCache+ reduces CCache’s average, p95, and p99 latencies by 26.9%, 14.7%, and 77.0%, respectively.
For the Alibaba workload, FMCache also outperforms NoCache. For example, at target throughput 0.15 MOPS, FMCache reduces NoCache’s average, p95 and p99 latencies by 66.1%, 65.4%, and 37.2%, respectively. For FMCache+ and CCache, at low throughput, they have comparable average and p95 latencies, while FMCache has slightly higher p99 latencies due to the switch’s cache maintenance overhead. At high throughput (e.g., 0.35 MOPS), FMCache+ reduces CCache’s average, p95, and p99 latencies by 25.5%, 63.7%, and 15.5%, respectively.
8.3. Impact of Workload Settings
We examine FMCache’s robustness across different workload settings. In the interest of space, we focus on Alibaba (with the largest write ratio) and Thumb (with the largest read ratio), while Training and LinkedIn show consistent behaviors with Thumb due to their similar read ratios.
(Exp#5) Impact of file access frequency assignment. We generate a sequence of access frequencies in descending order based on a power-law distribution with an exponent of 0.9, and another sequence of files based on different sorting orders. We assign the $i$-th access frequency to the $i$-th file. We consider three sorted file sequence: (i) high-level-first (HLF), which sorts files in descending order of their levels (i.e., files at higher levels have higher access frequencies), (ii) low-level-first (LLF), which sorts files in ascending order of their levels (i.e., files at lower levels have higher access frequencies), and (iii) random (our default), which generates a random sequence of files across all levels. Figure 11 shows that FMCache and FMCache+ still outperform NoCache and CCache, respectively, under different file access frequency assignments. For example, in Thumb (the most read-intensive), FMCache and FMCache+ increase the throughput of NoCache and CCache by 53.7-69.0% and 30.2-44.7%, respectively.
(Exp#6) Impact of access skewness. We vary the skewness level for access frequencies under the uniform distribution and varying power-law distributions with an exponent of 0.8, 0.9 (our default), and 1.0. A larger exponent implies a more skewed access pattern. Figure 12 shows that under the uniform workload, FMCache and FMCache+ have slightly less throughput than NoCache and CCache, respectively, by up to 5.0%, since most requests come from uncached paths and are served by servers, while FMCache and FMCache+ incur extra cache maintenance overhead.
Refer to caption Refer to caption
(a) Alibaba
(b) Thumb
Figure 11. (Exp#5) Impact of file access frequency assignment.
Refer to caption Refer to caption
(a) Alibaba
(b) Thumb
Figure 12. (Exp#6) Impact of access skewness.
Refer to caption Refer to caption
(a) Alibaba
(b) Thumb
Figure 13. (Exp#7) Impact of maximum path depth.
[…]
For Alibaba, as the exponent increases from 0.9 to 1.0, the throughput of FMCache and FMCache+ decreases. The reason is that Alibaba has the largest write ratio and incurs significant cache maintenance overhead. Nevertheless, at exponent 1.0, FMCache and FMCache+ still increase the throughput of NoCache and CCache by 37.6% and 25.5%, respectively.
(Exp#7) Impact of maximum path depth. We vary the maximum path depth as 3, 5, 7, and 9 (our default). Figure 13 shows that FMCache and FMCache+ always outperform NoCache and CCache, respectively. For example, in Thumb, FMCache and FMCache+ increase the throughput of NoCache and CCache by 53.7-77.9% and 34.1-36.3%, respectively.
(Exp#8) Impact of dynamic workloads. We consider dynamic workloads with varying access frequencies of files over time. We follow the prior studies (Li et al., 2016; Jin et al., 2017; Sheng et al., 2023) to generate the hot-in dynamic pattern, which periodically selects the 100 least-frequently accessed files, re-assigns them with the highest access frequencies, and adjusts the access frequencies of other files accordingly to maintain a power-law distribution. We set the change period as 20 seconds and run each scheme for 200 seconds to measure per-second throughput. We disable server rotation as the system states change under dynamic workloads; instead, we issue workloads to the two physical servers and measure performance directly.
Figure 14 shows that for Thumb, FMCache and FMCache+ show performance dips due to periodic changes of file access frequencies. Before new hot records are admitted, performance dips occur, but FMCache and FMCache+ quickly admit new hot records and return to high performance with path-aware cache management (§4). Also, local hash collision resolution (§6) incurs minimal overhead to cache admission and eviction. For Alibaba, which has the largest write ratio, FMCache and FMCache+ have marginal performance gains over NoCache and CCache, respectively. Nevertheless, FMCache and FMCache+ still effectively respond to dynamic workloads.
Refer to caption Refer to caption
(a) Alibaba
(b) Thumb
Figure 14. (Exp#8) Impact of dynamic workloads.
8.4. Switch Deployment
(Exp#9) Switch resource usage. We measure the resource usage on the Tofino switch for (i) SRAM (15 MiB), (ii) number of stages (12), (iii) number of ALUs (48), and (iv) PHV size (768 bytes). We also quote the resource usage of two in-switch key-value caches, NetCache (Jin et al., 2017) and FarReach (Sheng et al., 2023). Table 3 shows that NoCache and CCache use the least resources for L2/L3 forwarding, while FMCache and FMCache+ also support L2/L3 forwarding to process packets unrelated to metadata operations at line rate (3.2 Tbps for the Tofino switch (swi, [n. d.])). Additionally, FMCache and FMCache+ consume additional resources for caching (e.g., using SRAM and ALUs to cache metadata, track access frequencies, and maintain lock counters, and using PHVs to parse metadata requests), yet their resource usage is comparable to NetCache and FarReach (state-of-the-art in-switch key-value caches).
Scheme SRAM (KiB) # Stages # ALUs PHV size (bytes)
NoCache 288 (1.9%) 4 (33.3%) 0 (0%) 256 (33.3%)
CCache 288 (1.9%) 4 (33.3%) 0 (0%) 256 (33.3%)
NetCache (Jin et al., 2017) 8800 (57.3%) 12 (100%) 45 (93.8%) 528 (68.8%)
FarReach (Sheng et al., 2023) 8992 (58.5%) 12 (100%) 45 (93.8%) 499 (65.0%)
FMCache 8976 (58.4%) 12 (100%) 47 (97.6%) 712 (92.7%)
FMCache+ 8976 (58.4%) 12 (100%) 47 (97.6%) 712 (92.7%)
Table 3. (Exp#9) Switch resource usage (numbers in brackets refer to the fractions of used resources over total available ones). Numbers of NetCache and FarReach are quoted from (Sheng et al., 2023).
9. Related Work
Scaling file-system metadata management. Metadata partitioning distributes file-system management across multiple servers for scalability. There are two primary approaches: (i) dynamic sub-tree partitioning, which distributes namespace sub-trees across servers (Weil et al., 2006; Sevilla et al., 2015; Wang et al., 2021a), and (ii) hash-based partitioning, which distributes file-system metadata across servers via hashing (Patil and Gibson, 2011; Thomson and Abadi, 2015; Niazi et al., 2017; Pan et al., 2021; Liao and Abadi, 2023; Wang et al., 2023). FMCache complements these approaches as an in-switch cache that absorbs operations upstream of file-system metadata layers.
Client-side caching. PanFS (Welch et al., 2008) caches file and directory metadata and provides callbacks for cache consistency. IndexFS (Ren et al., 2014) and LocoFS (Li et al., 2017a) use lease-based caching for metadata management and invalidate cache entries upon lease expiration, but incur high overhead for renewing cache entries’ leases. InfiniFS (Lv et al., 2022) applies lazy invalidation for directory access metadata to limit the overhead of lease-based caching. Client-side caching often incurs high client-side complexity and overhead in maintaining cache consistency across a large number of clients. FMCache simplifies cache consistency by caching file-system metadata in a programmable switch that lies on the critical paths of multiple clients.
In-switch caching. Programmable switches have been extensively studied for concurrency control (Li et al., 2017b; Zhu et al., 2019), network monitoring (Gupta et al., 2018), replication coordination (Jin et al., 2018), remote procedure calls (Zhao et al., 2023), key-value stream aggregation (He et al., 2023), and distributed lock management (Yu et al., 2020; Zhang et al., 2024). Several studies explore in-switch caching. SwitchKV (Li et al., 2016) and NetCache (Jin et al., 2017) design write-through in-switch caching for read-intensive workloads, and DistCache (Liu et al., 2019) designs distributed write-through caching across multiple switches. Pegasus (Li et al., 2020) and TurboKV (Eldakiky et al., 2021) cache replica-to-server mappings for replica selection. Mind (Lee et al., 2021) caches object-to-memory mappings for disaggregated memory systems. Concordia (Wang et al., 2021b) tracks cache copy locations and states to address concurrency in shared memory systems. FarReach (Sheng et al., 2023) designs fault-tolerant write-back caching for write-intensive workloads. The above in-switch caches are designed for key-value stores, which differ from file-system semantics. AsyncFS (Xu et al., 2024) proposes in-switch tracking of directory updates, while maintaining a client-side metadata cache for path resolution. In contrast, FMCache moves file-system metadata caching to switches, and also complements client-side caching.
10. Conclusion
FMCache is an in-switch file-system metadata caching framework, aiming to achieve high throughput and load balancing for distributed file-system metadata services. It employs path-aware cache management, multi-level read-write locking, and local hash collision resolution. Experiments on a Tofino-switch testbed show that FMCache achieves significant throughput gains and complements client-side caching.

### v2

@@
-spacing=nonfrench
-Abstract.
+Fletch: File-System Metadata Caching in Programmable Switches
+Abstract
 Fast and scalable metadata management across multiple metadata servers is crucial for distributed file systems to handle numerous files and directories.
 Client-side caching of frequently accessed metadata can mitigate server loads, but incurs significant overhead and complexity in maintaining cache consistency when the number of clients increases.
-We propose FMCache, an in-switch file-system metadata caching framework that leverages programmable switches to serve file-system metadata requests from multiple clients directly in the switch data plane.
-Unlike prior in-switch key-value caching approaches, FMCache addresses file-system-specific path dependencies under stringent switch resource constraints.
-We implement FMCache atop Hadoop HDFS and evaluate it on a Tofino-switch testbed using real-world file-system metadata workloads.
-FMCache achieves up to 181.6% higher throughput than vanilla HDFS and complements client-side caching with additional throughput gains of up to 139.6%.
+We explore caching in programmable switches by serving file-system metadata requests from multiple clients on the switch data plane.
+Despite prior efforts on in-switch key-value caching, they fail to address the path dependencies specific to file-system semantics.
+We propose Fletch, an in-switch file-system metadata caching framework that leverages programmable switches to serve file-system metadata requests from multiple clients directly in the switch data plane.
+Unlike prior in-switch key-value caching approaches, Fletch addresses file-system-specific path dependencies under stringent switch resource constraints.
+We implement Fletch atop Hadoop HDFS and evaluate it on a Tofino-switch testbed using real-world file-system metadata workloads.
+Fletch achieves up to 181.6% higher throughput than vanilla HDFS and complements client-side caching with additional throughput gains of up to 139.6%.
 It also incurs low latencies and limited switch resource usage.
-1. Introduction
+I Introduction
 Scaling file-system metadata management across multiple metadata servers is crucial for distributed file systems to handle billions of files and directories.
-Field studies, both classical (Roselli et al., 2000; Harter et al., 2012) and recent (Wang et al., 2023), indicate that metadata operations dominate file-system requests.
-For example, 67-96% of file-system requests in Baidu AI Cloud are metadata-related (Wang et al., 2023).
-As the number of files and directories grows, especially in workloads dominated by small files (Carns et al., 2009; Harter et al., 2014), metadata operations, such as inode lookups, permission checks, and directory traversals, become performance bottlenecks.
+Field studies, both classical [roselli00, harter12] and recent [wang23], indicate that metadata operations dominate file-system requests.
+For example, 67-96% of file-system requests in Baidu AI Cloud are metadata-related [wang23].
+As the number of files and directories grows, especially in workloads dominated by small files [carns09, harter14], metadata operations, such as inode lookups, permission checks, and directory traversals, become performance bottlenecks.
 Caching frequently accessed metadata on the client side is a common strategy to mitigate loads on metadata servers.
-File-system metadata access patterns in practice tend to be skewed, with a small fraction of files and directories being accessed far more frequently than others (Leung et al., 2008; Abad et al., 2012; Ananthanarayanan et al., 2011); for example, less than 3% of files account for 34-39% of requests in Yahoo’s HDFS clusters (Abad et al., 2012).
+File-system metadata access patterns in practice tend to be skewed, with a small fraction of files and directories being accessed far more frequently than others [leung08, abad12, ananthanarayanan11]; for example, less than 3% of files account for 34-39% of requests in Yahoo’s HDFS clusters [abad12].
 However, client-side caching incurs significant overhead in maintaining cache consistency across numerous clients, as metadata updates require notifications to all clients to invalidate cached metadata.
-Existing client-side caching approaches (Welch et al., 2008; Ren et al., 2014; Li et al., 2017a; Lv et al., 2022) cache limited metadata information to mitigate path resolution overhead (§2.1) but rely on metadata servers for accessing metadata contents, so the performance gains are limited.
-Programmable switches (Bosshart et al., 2013) offer a promising alternative by enabling in-switch caching across multiple clients and leveraging a centralized view of requests, so as to eliminate switch-to-server transmissions and server-side processing.
-While prior work has explored in-switch caching for key-value stores (Li et al., 2016; Jin et al., 2017; Liu et al., 2019; Sheng et al., 2023; Li et al., 2020; Eldakiky et al., 2021; Lee et al., 2021), in-switch caching for file-system metadata poses unique challenges not directly addressed before.
+Existing client-side caching approaches [welch08, ren14, li17, lv22] cache limited metadata information to mitigate path resolution overhead (§II-A) but rely on metadata servers for accessing metadata contents, so the performance gains are limited.
+Programmable switches [bosshart13] offer a promising alternative by enabling in-switch caching across multiple clients and leveraging a centralized view of requests, so as to eliminate switch-to-server transmissions and server-side processing.
+While prior work has explored in-switch caching for key-value stores [li16, jin17, liu19, sheng25, li20, eldakiky20, lee21], in-switch caching for file-system metadata poses unique challenges not directly addressed before.
 First, file-system pathnames, unlike fixed-length keys, have large and variable sizes and cannot readily fit into limited switch resources.
@@
 Third, cache lookups for the metadata of files and directories require multiple iterations under the strict switch programming model, thereby complicating concurrent cache updates and lookups while maintaining consistency and performance.
-We propose FMCache, an in-switch file-system metadata caching framework tailored for distributed file systems under skewed, read-intensive workloads.
-FMCache extends in-switch key-value caches by specifically addressing file-system path dependencies through techniques with performance and correctness guarantees: (i) path-aware cache management to account for path dependencies in cache admission and eviction; (ii) multi-level read-write locking to enable high-performance concurrent cache lookups and updates; and (iii) local hash collision resolution to efficiently and correctly map variable-length paths into fixed-length keys.
-We implemented FMCache in P4 (Bosshart et al., 2014) and compiled it into the Tofino switch chipset (tof, [n.
-d.]; swi, [n.
-d.]).
-We integrated FMCache with Hadoop HDFS (hdf, [n.
-d.]a), while preserving HDFS semantics.
-We also implemented a state-of-the-art client-side caching approach (Ren et al., 2014; Lv et al., 2022).
-Our evaluation on four real-world workloads (Ren et al., 2014; Lv et al., 2022; Xu et al., 2024) shows that under 128 simulated metadata servers, FMCache achieves up to 181.6% higher throughput than vanilla HDFS, while client-side caching coupled with FMCache achieves up to 139.6% higher throughput than without FMCache.
-2. Background and Motivation
-2.1. File-System Metadata Management
+We propose Fletch, an in-switch file-system metadata caching framework tailored for distributed file systems under skewed, read-intensive workloads.
+Fletch extends in-switch key-value caches by specifically addressing file-system path dependencies through techniques with performance and correctness guarantees: (i) path-aware cache management to account for path dependencies in cache admission and eviction; (ii) multi-level read-write locking to enable high-performance concurrent cache lookups and updates; and (iii) local hash collision resolution to efficiently and correctly map variable-length paths into fixed-length keys.
+We implemented Fletch in P4 [bosshart14] and compiled it into the Tofino switch chipset [tofino, switchthpt].
+We integrated Fletch with Hadoop HDFS [hdfs], while preserving HDFS semantics.
+We also implemented a state-of-the-art client-side caching approach [ren14, lv22].
+Our evaluation on four real-world workloads [ren14, lv22, xu24] shows that under 128 simulated metadata servers, Fletch achieves up to 181.6% higher throughput than vanilla HDFS, while client-side caching coupled with Fletch achieves up to 139.6% higher throughput than without Fletch.
+We release the source code of Fletch at https://github.com/calanquee/Fletch.
+II Background and Motivation
+II-A File-System Metadata Management
 File systems organize data in a hierarchical, tree-based namespace, where files and directories are managed as leaf and non-leaf nodes, respectively.
@@
 We say that $p$ is the parent of $q$, and $q$ is a child of $p$, if $p$ is an ancestor of $q$ with exactly one less level than $q$ (e.g., /a is the parent of /a/b, and /a/b is a child of /a).
-In this work, we focus on Hadoop Distributed File System (HDFS) (Shvachko et al., 2010), where a namenode manages the namespace and metadata.
+In this work, we focus on Hadoop Distributed File System (HDFS) [shvachko10], where a namenode (i.e., metadata server) manages the namespace and metadata.
 HDFS supports various metadata operations for files (e.g., create, delete, open, and close), directories (e.g., mkdir and rmdir), attributes (e.g., chmod, chown, and utime), and data organization (i.e., rename, readdir, and stat).
 Metadata operations rely on path resolution, which parses and traverses each level of a path to verify metadata for existence and permissions.
-For scalability, HDFS employs Router-Based Federation (RBF) (hdf, [n.
-d.]b) to distribute namespace and metadata management across multiple namenodes.
-Figure 1. Data plane of a programmable switch.
-2.2. Programmable Switches
-Programmable switches (Bosshart et al., 2013) enable customized packet processing tasks beyond traditional packet forwarding.
-We focus on Tofino switches (tof, [n.
-d.]), which are programmed using the P4 language (Bosshart et al., 2014) and built on the Protocol-Independent Switch Architecture (PISA) based on the Reconfigurable Match-Action Table (RMT) paradigm (Bosshart et al., 2013).
-We target the generic RMT architecture without relying on any Tofino-specific features, so our design is extensible to other RMT-based switch platforms (bro, [n.
-d.]; Chole et al., 2017; hua, [n.
-d.]).
+For scalability, HDFS employs Router-Based Federation (RBF) [hdfsrbf] to distribute namespace and metadata management across multiple namenodes.
+Figure 1: Data plane of a programmable switch.
+II-B Programmable Switches
+Programmable switches [bosshart13] enable customized packet processing tasks beyond traditional packet forwarding.
+We focus on Tofino switches [tofino], which are programmed using the P4 language [bosshart14] and built on the Protocol-Independent Switch Architecture (PISA) based on the Reconfigurable Match-Action Table (RMT) paradigm [bosshart13].
+We target the generic RMT architecture without relying on any Tofino-specific features, so our design is extensible to other RMT-based switch platforms [broadcomswitch, chole17, huaweiswitch].
 A programmable switch comprises a data plane and a control plane.
@@
 Recirculation should be cautiously used, as it consumes extra switch resources and slows down packet forwarding.
-In this work, we leverage recirculation for read requests that require in-switch path resolution (§4.1) and write requests that await locks (§5.2).
-2.3. Challenges
+In this work, we leverage recirculation for read requests that require in-switch path resolution (§IV-A) and write requests that await locks (§V-B).
+II-C Challenges
 Designing in-switch metadata caching for distributed file systems poses the following challenges.
@@
 Switches have limited resources and complicate caching implementation.
-For example, Tofino switches (tof, [n.
-d.]) provide only 12 stages per ingress or egress pipeline, each processing up to 16 bytes due to ALU word size restrictions.
-The PHV size is capped (e.g., 768 bytes in Tofino switches (tof, [n.
-d.])), and all pipelines share limited SRAM.
-For comparisons, NetCache (Jin et al., 2017) allocates SRAM for access frequency tracking and supports only a maximum value size of 128 bytes for key-value records.
+For example, Tofino switches [tofino] provide only 12 stages per ingress or egress pipeline, each processing up to 16 bytes due to ALU word size restrictions.
+The PHV size is capped (e.g., 768 bytes in Tofino switches [tofino]), and all pipelines share limited SRAM.
+For comparison, NetCache [jin17] allocates SRAM for access frequency tracking and supports only a maximum value size of 128 bytes for key-value records.
 File-system metadata caching is particularly resource-intensive.
-HDFS file or directory names can reach 255 bytes (def, [n.
-d.]) and aggravate SRAM and PHV overhead.
-Path resolution for a path (e.g., /a/b.txt) requires accessing metadata for its ancestors (i.e., / and /a) (§2.1).
+HDFS file or directory names can reach 255 bytes [defaultsetting] and aggravate SRAM and PHV overhead.
+Path resolution for a path (e.g., /a/b.txt) requires accessing metadata for its ancestors (i.e., / and /a) (§II-A).
 Caching metadata for all paths and their ancestors is desirable but challenging given limited SRAM.
 Splitting path processing across multiple levels requires careful synchronization across stages.
-To mitigate switch resource usage, existing in-switch key-value caches (Jin et al., 2017; Liu et al., 2019; Sheng et al., 2023) offload cache admission and eviction to a centralized controller, but excessive switch-to-controller communications incur high latencies.
+To mitigate switch resource usage, existing in-switch key-value caches [jin17, liu19, sheng25] offload cache admission and eviction to a centralized controller, but excessive switch-to-controller communications incur high latencies.
 Challenge 2: Cache consistency.
@@
 Thus, the read request can access a mix of pre- and post-updated metadata, leading to inconsistencies.
-3. Design Overview
-3.1. Goals and Assumptions
-FMCache is an in-switch metadata caching framework for distributed file systems, aiming to achieve high throughput and low latencies for file-system metadata requests across multiple metadata servers.
-It is currently designed for HDFS (§2.1), supporting all HDFS metadata operations via its interface (§7) and preserving HDFS semantics.
+III Design Overview
+III-A Goals and Assumptions
+Fletch is an in-switch metadata caching framework for distributed file systems, aiming to achieve high throughput and low latencies for file-system metadata requests across multiple metadata servers.
+It is currently designed for HDFS (§II-A), supporting all HDFS metadata operations via its interface (§VII) and preserving HDFS semantics.
 It is also extensible to other distributed file systems via their respective APIs.
-FMCache targets read-intensive file-system metadata workloads, as observed in prior studies (Ren et al., 2014; Lv et al., 2022; Xu et al., 2024).
-For example, in a LinkedIn HDFS cluster, 84% of 145 million metadata operations are lookups, with only 9% creates and 7% updates (Ren et al., 2014); in Alibaba’s Pangu file system, over 60% of metadata operations are reads (Lv et al., 2022).
-To ensure consistency, FMCache adopts write-through caching (as in prior in-switch key-value caches (Li et al., 2016; Liu et al., 2017; Jin et al., 2017; Liu et al., 2019)) to update both the in-switch cache and server-side metadata in writes before acknowledging clients.
-Also, since path resolution always starts at the root, FMCache assumes the root directory’s metadata is persistently cached, with its permissions unchanged and the root not being deleted (Niazi et al., 2017).
-3.2. Architecture and Design Roadmap
-Figure 2. FMCache’s architecture.
-Figure 2 depicts FMCache’s architecture, which comprises both control-plane and data-plane components.
+Fletch targets read-intensive file-system metadata workloads, as observed in prior studies [ren14, lv22, xu24].
+For example, in a LinkedIn HDFS cluster, 84% of 145 million metadata operations are lookups, with only 9% creates and 7% updates [ren14]; in Alibaba’s Pangu file system, over 60% of metadata operations are reads [lv22].
+To ensure consistency, Fletch adopts write-through caching (as in prior in-switch key-value caches [li16, liu17, jin17, liu19]) to update both the in-switch cache and server-side metadata in writes before acknowledging clients.
+With write-through caching, Fletch avoids possible data loss caused by switch failures as metadata servers always possess the latest file-system metadata.
+Also, since path resolution always starts at the root, Fletch assumes the root directory’s metadata is persistently cached, with its permissions unchanged and the root not being deleted [niazi17].
+III-B Architecture and Design Roadmap
+Figure 2: Fletch’s architecture.
+Figure 2 depicts Fletch’s architecture, which comprises both control-plane and data-plane components.
 In the control plane, a controller manages in-switch cache admission and eviction, while in the data plane, multiple clients send (receive) file-system metadata requests (responses) via the switch to (from) multiple metadata servers (or servers for short).
-FMCache employs three techniques to enable efficient in-switch file-system metadata caching.
-Path-aware cache management (§4).
-FMCache caches file-system metadata in the switch data plane.
-FMCache hashes variable-length paths into fixed-length keys and manages file-system metadata in key-value pairs as in existing in-switch key-value caches (Li et al., 2016; Jin et al., 2017; Liu et al., 2019; Sheng et al., 2023).
-But unlike in-switch key-value caches, FMCache addresses path dependencies across hierarchical levels during cache admission and eviction.
-Multi-level read-write locking (§5).
-To synchronize concurrent metadata operations, FMCache adopts multi-level read-write locking using (i) lock counters as read-write locks for different path levels and (ii) validation flags for verifying the validity of cached paths.
+Fletch employs three techniques to enable efficient in-switch file-system metadata caching.
+Path-aware cache management (§IV).
+Fletch caches file-system metadata in the switch data plane.
+Fletch hashes variable-length paths into fixed-length keys and manages file-system metadata in key-value pairs as in existing in-switch key-value caches [li16, jin17, liu19, sheng25].
+But unlike in-switch key-value caches, Fletch addresses path dependencies across hierarchical levels during cache admission and eviction.
+Multi-level read-write locking (§V).
+To synchronize concurrent metadata operations, Fletch adopts multi-level read-write locking using (i) lock counters as read-write locks for different path levels and (ii) validation flags for verifying the validity of cached paths.
 It ensures reliable locking and unlocking during path resolution even in unreliable networks.
-Local hash collision resolution (§6).
+Local hash collision resolution (§VI).
 Mapping paths to fixed-size keys in the switch data plane can cause hash collisions, leading to incorrect metadata retrieval.
 For example, if two paths /a and /b share the same hash and /a is cached, a request for /b can erroneously return /a’s metadata.
-FMCache proposes a local hash collision resolution approach, where the controller assigns unique values, called tokens, to hash-colliding paths and distributes the tokens to clients, servers, and the switch, so as to allow local resolution of hash collisions without the controller’s intervention for every request.
-FMCache ensures no incorrect metadata retrieval.
-4. Path-Aware Cache Management
-FMCache incorporates path awareness into cache admission and eviction by caching frequently accessed paths and their ancestors, so as to mitigate cache misses during path resolution.
+Fletch proposes a local hash collision resolution approach, where the controller assigns unique values, called tokens, to hash-colliding paths and distributes the tokens to clients, servers, and the switch, so as to allow local resolution of hash collisions without the controller’s intervention for every request.
+Fletch ensures no incorrect metadata retrieval.
+IV Path-Aware Cache Management
+Fletch incorporates path awareness into cache admission and eviction by caching frequently accessed paths and their ancestors, so as to mitigate cache misses during path resolution.
 This ensures that if a path is cached, its ancestors are also cached, thereby allowing the reuse of cached metadata across different paths with the common ancestors.
-4.1. Path Representation
-FMCache maps a file-system pathname into fixed-size hash keys for efficient switch operations.
-For a read operation, a FMCache client partitions a path into multiple levels (including the root) and computes a hash for each level.
+IV-A Path Representation
+Fletch maps a file-system pathname into fixed-size hash keys for efficient switch operations.
+For a read operation, a Fletch client partitions a path into multiple levels (including the root) and computes a hash for each level.
 For example, /a/b/c.txt is partitioned into /, /a, /a/b, and /a/b/c.txt, with each level being hashed.
 To avoid redundant computations, the hash of the root directory / is pre-computed and cached on the client side.
-For a write operation, the client computes a hash for the complete path without partitioning, as all writes are forwarded to the server (that holds the namespace) for path resolution under write-through caching (§3.1).
-Currently, FMCache uses the first 64 bits of a 128-bit MD5 hash as the hash value.
-FMCache caches file-system metadata as key-value records in the switch.
+For a write operation, the client computes a hash for the complete path without partitioning, as all writes are forwarded to the server (that holds the namespace) for path resolution under write-through caching (§III-A).
+Currently, Fletch uses the first 64 bits of a 128-bit MD5 hash as the hash value; note that MD5 is used for fast non-cryptographic hashing; collision resistance is addressed separately via tokens (§VI).
+Fletch caches file-system metadata as key-value records in the switch.
 Each record is identified by its hash key, and its value contains file or directory metadata.
@@
 Each file’s metadata has 40 bytes, and each directory’s has 24 bytes.
-FMCache manages cache across multiple stages in the switch.
-It leverages MATs across two stages in a single ingress pipeline to perform cache lookups based on hash keys, so as to avoid cross-pipeline synchronization (§5) and enforce logical dependency (i.e., cache hit/miss status is resolved before metadata retrieval can proceed).
+Fletch manages cache across multiple stages in the switch.
+It leverages MATs across two stages in a single ingress pipeline to perform cache lookups based on hash keys, so as to avoid cross-pipeline synchronization (§V) and enforce logical dependency (i.e., cache hit/miss status is resolved before metadata retrieval can proceed).
 Since packets are strictly processed in ingress-to-egress order, this placement enables fast cache status checks.
-The switch manages metadata in 32 registers across eight stages in the egress pipelines (§7).
+The switch manages metadata in 32 registers across eight stages in the egress pipelines (§VII).
 To simplify routing, the switch caches metadata in the egress pipeline that is connected to the corresponding server.
-Clients issue metadata requests, including hash keys and full pathnames, to the switch, which reports frequently accessed pathnames to the controller during cache admission (§4.2).
+Clients issue metadata requests, including hash keys and full pathnames, to the switch, which reports frequently accessed pathnames to the controller during cache admission (§IV-B).
 For reads, the switch performs path resolution by issuing cache lookups and permission checks for each level, starting from the root, and recirculates the request to the ingress pipeline for each next level.
@@
 For a cache miss, the switch forwards the request to the server.
-Due to limited stages and memory access constraints (§2.2), FMCache cannot perform path resolution for multiple path levels in a single pass, but instead leverages recirculation for path resolution while limiting overhead (§8.2).
-4.2. Cache Admission and Eviction Workflows
+Due to limited stages and memory access constraints (§II-B), Fletch cannot perform path resolution for multiple path levels in a single pass, but instead leverages recirculation for path resolution while limiting overhead (§VIII-B).
+IV-B Cache Admission and Eviction Workflows
 Data structures.
-FMCache monitors path access frequencies for read operations (excluding access to ancestors during path resolution).
-For uncached paths, which dominate path access traffic, FMCache uses a Count-Min Sketch (CMS) (Cormode and Muthukrishnan, 2005), as in prior studies (Jin et al., 2017; Liu et al., 2019; Sheng et al., 2023), to estimate their access frequencies within fixed-size memory with provable error bounds.
-For cached paths, FMCache tracks their exact access frequencies using a frequency counter array.
+Fletch monitors path access frequencies for read operations (excluding access to ancestors during path resolution).
+For uncached paths, which dominate path access traffic, Fletch uses a Count-Min Sketch (CMS) [cormode05], as in prior studies [jin17, liu19, sheng25], to estimate their access frequencies within fixed-size memory with provable error bounds.
+For cached paths, Fletch tracks their exact access frequencies using a frequency counter array.
 The switch periodically reports the access frequencies of all cached paths to the controller for eviction decisions, and resets the CMS and frequency counter array after each reporting.
 Cache admission.
-FMCache’s cache admission is triggered by the switch data plane.
+Fletch’s cache admission is triggered by the switch data plane.
 During runtime, the switch monitors the access frequencies of uncached paths via the CMS and identifies hot paths (i.e., those exceeding a pre-defined CMS threshold) for admission.
 When a hot path $p$ is detected, the switch notifies the controller, which retrieves the metadata for $p$ and its uncached ancestors from the servers.
-The controller communicates with servers using UDP (and issues retransmissions if needed), bypassing the switch data plane to avoid critical-path overhead, as in prior work (Jin et al., 2017; Sheng et al., 2023).
+The controller communicates with servers using UDP (and issues retransmissions if needed), bypassing the switch data plane to avoid critical-path overhead, as in prior work [jin17, sheng25].
 The controller verifies cache capacity.
 If the cache is not full, the controller sends the hash keys and metadata for $p$ and its uncached ancestors to the switch for admission (note that the controller maintains a global view of all cached paths).
-Unlike NetCache (Jin et al., 2017) (without path awareness), FMCache admits the metadata for both $p$ and its uncached ancestors to ensure path-aware caching.
+Unlike NetCache [jin17] (without path awareness), Fletch admits the metadata for both $p$ and its uncached ancestors to ensure path-aware caching.
+To guarantee cache coherence during cache admission, while the controller is adding a path to the cache, write requests to this path are blocked at the server until the server is notified that the cache admission is finished [jin17].
+Since file-system metadata workloads are read-intensive (§III-A), the blocking action has limited impact on the overall performance in general.
 Cache eviction.
@@
 The controller recursively includes ancestors as candidates until reaching an ancestor with multiple cached children or the root.
-The rationale of this recursive inclusion is that the selected ancestors are unlikely accessed by themselves, as directory operations are infrequent in practice (Ren et al., 2014; Lv et al., 2022; Xu et al., 2024) (e.g., only 4.2% of metadata operations in Alibaba’s Pangu file system are directory-related (Lv et al., 2022)).
+The rationale of this recursive inclusion is that the selected ancestors are unlikely accessed by themselves, as directory operations are infrequent in practice [ren14, lv22, xu24] (e.g., only 4.2% of metadata operations in Alibaba’s Pangu file system are directory-related [lv22]).
 The controller selects candidates up to a pre-defined threshold, currently set as twice the number of paths to be admitted (i.e., $p$ and its uncached ancestors), as the periodically reported access frequencies may vary over time and selecting more candidates than being admitted can avoid mistakenly evicting hot paths.
@@
 Finally, the controller notifies the switch to evict the selected paths and admits $p$ and its uncached ancestors.
-Figure 3. Example of cache admission and eviction workflows.
+Figure 3: Example of cache admission and eviction workflows.
 Example.
@@
 The controller notifies the switch to evict /a and /a/b.txt, and admits /c and /c/d.txt.
-5. In-Switch Read-Write Locking
-FMCache adopts multi-level read-write locking to ensure cache consistency under concurrent metadata updates.
+V In-Switch Read-Write Locking
+Fletch adopts multi-level read-write locking to ensure cache consistency under concurrent metadata updates.
 The locking mechanism operates entirely within the switch data plane to avoid controller overhead.
-5.1. Lock Design
-FMCache employs multiple lock counter arrays and a validation array, both implemented as register arrays in the switch data plane and initialized with zero entries.
+V-A Lock Design
+Fletch employs multiple lock counter arrays and a validation array, both implemented as register arrays in the switch data plane and initialized with zero entries.
 Each cached path is associated with one lock counter (a slot in a lock counter array) and one validation flag (a slot in the validation array), indexed by the path’s hash key.
@@
 Each of its lock counters records the number of active read requests for a cached path at that level.
-Due to switch resource constraints, FMCache allocates eight counter arrays of 65,536 16-bit counters each, while each counter supports up to 65,535 concurrent read requests.
-The first seven arrays are assigned to levels 1 through 7 (e.g., /a for level 1 (§2.1)), while the eighth array handles all remaining deeper levels (using the hash key of level 8).
-This design maximizes concurrent access to shallow paths, based on empirical evidence that most metadata requests are aggregated at small depths (Agrawal et al., 2007; Meyer and Bolosky, 2012; Douceur and Bolosky, 1999); for example, 90% (Agrawal et al., 2007; Meyer and Bolosky, 2012), or even close to 100% (Douceur and Bolosky, 1999), of accessed paths have a depth of no more than 10.
-FMCache maps a cached path to a lock counter array based on its level, and uses its hash key’s last 16 bits to associate the path with a specific lock counter.
+Due to switch resource constraints, Fletch allocates eight counter arrays of 65,536 16-bit counters each, while each counter supports up to 65,535 concurrent read requests.
+The first seven arrays are assigned to levels 1 through 7 (e.g., /a for level 1 (§II-A)), while the eighth array handles all remaining deeper levels (using the hash key of level 8).
+This design maximizes concurrent access to shallow paths, based on empirical evidence that most metadata requests are aggregated at small depths [agrawal07, meyer12, douceur99]; for example, 90% [agrawal07, meyer12], or even close to 100% [douceur99], of accessed paths have a depth of no more than 10.
+Fletch maps a cached path to a lock counter array based on its level, and uses its hash key’s last 16 bits to associate the path with a specific lock counter.
 Validation array.
-The validation array (Jin et al., 2017; Sheng et al., 2023) tracks metadata validity for all cached paths.
+The validation array [jin17, sheng25] tracks metadata validity for all cached paths.
 A validation flag of one indicates valid metadata and allows reads from the cache, while a flag of zero indicates invalid or updating metadata and all reads are directed to servers.
 After a cache update, its validation flag is set to one to permit subsequent reads.
-5.2. Read and Write Flows
-FMCache places the validation array across all egress pipelines, co-located with value register arrays (§4.1) for efficient validity checks.
+V-B Read and Write Flows
+Fletch places the validation array across all egress pipelines, co-located with value register arrays (§IV-A) for efficient validity checks.
 Lock counter arrays reside in a single ingress pipeline to avoid cross-pipeline synchronization.
-FMCache redirects requests arriving at other ingress pipelines to the ingress pipeline holding lock counter arrays via cross-pipeline recirculation (§7).
+Fletch redirects requests arriving at other ingress pipelines to the ingress pipeline holding lock counter arrays via cross-pipeline recirculation (§VII).
 This placement is critical, as lock counter arrays should be positioned before the validation array to ensure correctness, and they cannot be placed in egress pipelines due to insufficient switch resources.
-In FMCache’s deployment, the switch data plane is not a bottleneck, and consistent packet processing ordering is maintained across ingress and egress pipelines, so FMCache ensures efficient and correct locking and validation operations.
+In Fletch’s deployment, the switch data plane is not a bottleneck, and consistent packet processing ordering is maintained across ingress and egress pipelines, so Fletch ensures efficient and correct locking and validation operations.
 Reads.
@@
 Single-path reads are served from the in-switch cache, while multi-path reads are forwarded to servers to ensure correctness, as descendant paths may be uncached and servers maintain the authoritative namespace to resolve partially caching scenarios.
-Multi-path reads are rare in practice (e.g., only 3.9% in Alibaba’s workloads (Lv et al., 2022); see Table 1).
-Upon receiving a (single-path) read request, the switch checks if the path’s last level is cached, implying that all its ancestors are also cached (§4).
+Multi-path reads are rare in practice (e.g., only 3.9% in Alibaba’s workloads [lv22]; see Table I).
+Upon receiving a (single-path) read request, the switch checks if the path’s last level is cached, implying that all its ancestors are also cached (§IV).
 If so, the switch increments the lock counter for each path level by one and resolves the path via recirculation.
@@
 Any ACK loss can cause duplicate lock decrements via retransmissions and violate correctness.
-FMCache handles ACK loss via a server-switch sequence-number protocol.
+Fletch handles ACK loss via a server-switch sequence-number protocol.
 Each server tags a lock-related response with a local sequence number (initially 0), and increments the sequence number upon receiving an ACK from the switch.
@@
 Upon receiving a server’s response, the switch compares the embedded sequence number with its expected value: if they match, the switch increments the expected value (for the next response), processes lock updates, forwards the response to the client, and returns an ACK to the server; if the embedded sequence number is lower, it indicates a duplicate, so the switch only sends an ACK to the server to suppress further retransmissions.
-Figure 4. Example of processing a read request under multi-level read-write locking.
+Figure 4: Example of processing a read request under multi-level read-write locking.
 For example, consider three cached paths /a, /a/b, and /a/b/c.txt.
@@
 However, the short recirculation time makes this unlikely.
-FMCache’s locking ensures correctness and allows the write to proceed after sufficient recirculations.
+Fletch’s locking ensures correctness and allows the write to proceed after sufficient recirculations.
 (ii) Multi-path writes.
@@
 If the write is successfully completed, the server updates the cache for all cached descendants before the requested path, so that the requested path remains invalidated until all cached descendants are fully updated.
-By performing path resolution in a top-down manner, FMCache prevents reads from accessing cached descendants before the requested path is updated (i.e., until the cache updates for a multi-path write are completed), so as to ensure cache consistency.
-6. Local Hash Collision Resolution
-Figure 5. Example of token generation and distribution. Note that the controller also assigns tokens for $p$’s uncached ancestors in Steps 2-4. We omit the details for brevity.
-FMCache maps file-system pathnames to 64-bit hash keys (§4.1), so hash collisions are possible (albeit unlikely) and lead to incorrect metadata retrieval.
-While the controller holds a global view of all cached paths (§4.2), querying the controller to resolve hash collisions for every request is impractical due to high switch-to-controller latencies (§2.3).
-FMCache adopts a local, token-based hash collision resolution mechanism by synchronizing the controller’s global view of cached paths with the switch, clients, and servers, so as to ensure correctness without compromising performance.
+By performing path resolution in a top-down manner, Fletch prevents reads from accessing cached descendants before the requested path is updated (i.e., until the cache updates for a multi-path write are completed), so as to ensure cache consistency.
+VI Local Hash Collision Resolution
+Figure 5: Example of token generation and distribution. Note that the controller also assigns tokens for $p$’s uncached ancestors in Steps 2-4. We omit the details for brevity.
+Fletch maps file-system pathnames to 64-bit hash keys (§IV-A), so hash collisions are possible (albeit unlikely) and lead to incorrect metadata retrieval.
+While the controller holds a global view of all cached paths (§IV-B), querying the controller to resolve hash collisions for every request is impractical due to high switch-to-controller latencies (§II-C).
+Fletch adopts a local, token-based hash collision resolution mechanism by synchronizing the controller’s global view of cached paths with the switch, clients, and servers, so as to ensure correctness without compromising performance.
 Hash collision resolution with tokens.
@@
 Thus, we can associate valid tokens with up to 255 paths with the same colliding hash key.
-The probability of two paths colliding is $1/{2}^{32}$ based on the birthday paradox (Flajolet et al., 1992), making it unlikely for more than 255 paths to share the same hash key in a large namespace.
+For 64-bit hash keys, the collision probability between any two paths is $1/2^{64}$, making it unlikely for more than 255 paths to share the same hash key in a large namespace.
 Thus, 8-bit tokens suffice for hash collision resolution.
-FMCache maintains two unordered map structures: (i) a path-token map, which records paths and their tokens (path-token pairs), and (ii) a hash-token map, which records hash keys and their tokens (hash-token pairs).
+Fletch maintains two unordered map structures: (i) a path-token map, which records paths and their tokens (path-token pairs), and (ii) a hash-token map, which records hash keys and their tokens (hash-token pairs).
 The controller keeps both maps, each client and server holds a path-token map, and the switch holds a hash-token map.
@@
 In the worst case, the controller may store entries for all paths ever cached.
-For scalability, the controller can use persistent key-value stores (e.g., RocksDB (roc, [n.
-d.])) for the map structures.
+For scalability, the controller can use persistent key-value stores (e.g., RocksDB [rocksdb]) for the map structures.
 The controller distributes each admitted path and its token to the switch and the relevant server that holds the path ( 4).
@@
 These entries are removed during cache eviction, as notified by the controller.
-Figure 6.
-Example of how a client updates its path-token map.
+Figure 6: Example of how a client updates its path-token map.
 Note that the client also attaches tokens of $p$’s ancestors in Step 1, the server returns the tokens for $p$’s ancestors in Step 5, and the client adds the tokens in Step 6.
@@
 Discussion.
-FMCache’s token-based mechanism ensures correctness with modest client complexity and overhead.
+Fletch’s token-based mechanism ensures correctness with modest client complexity and overhead.
 Clients must maintain path-token maps and attach tokens to requests, and incur a cache miss on the first read of a cached path due to the initial lack of a valid token.
@@
 Theoretically, the number of tokens grows linearly with the depth of a path.
-Nevertheless, as most paths have small depths (§5.1), the communication overhead is minimal.
-7. Implementation
-We have built a prototype of FMCache, including the controller, in-switch cache, clients, and servers.
+Nevertheless, as most paths have small depths (§V-A), the communication overhead is minimal.
+If the controller fails, Fletch needs to reboot the switch, controller, and servers to ensure correct cache retrieval.
+The current caches can be quickly recovered by exporting all caches before the reboot.
+VII Implementation
+We have built a prototype of Fletch, including the controller, in-switch cache, clients, and servers.
 Controller.
 The controller is implemented in C++ with 2.8 K LoC.
-It leverages APIs provided by the Tofino switch compiler (tof, [n.
-d.]) to interact with the switch data plane, including configurations and updates of MATs and registers, for cache admission and eviction (§4) and token management (§6).
+It leverages APIs provided by the Tofino switch compiler [tofino] to interact with the switch data plane, including configurations and updates of MATs and registers, for cache admission and eviction (§IV) and token management (§VI).
 In-switch cache.
-The in-switch cache is implemented in P4 with 5.1 K LoC and compiled for a Tofino switch (tof, [n.
-d.]; swi, [n.
-d.]).
+The in-switch cache is implemented in P4 with 5.1 K LoC and compiled for a Tofino switch [tofino, switchthpt].
 It comprises several components: (i) a hash-token map using MATs for cache lookups and local hash collision resolution, where each entry is a 9-byte key (8-byte hash key and 1-byte token); (ii) 32 register arrays of 32-bit slots for storing cache entries; (iii) a three-row CMS with three register arrays, each with 64K 16-bit slots; (iv) a frequency counter array with a register array of 32-bit slots; (v) lock counter arrays with eight register arrays, each with 16K 16-bit slots; (vi) a validation array with a register array of 1-bit slots; and (vii) a sequence counter array with a register array of 8-bit slots.
@@
 The C++ client driver supports multi-threaded workload execution, where each thread simulates a logical client.
-Each client maintains a path-token map for local hash collision resolution (§6).
+Each client maintains a path-token map for local hash collision resolution (§VI).
 It communicates with the switch via UDP-encapsulated packets customized for metadata requests, with re-transmission support for reliability.
-We integrate the driver with the mdtest (mdt, [n.
-d.]) benchmarking tool for HDFS metadata operations.
+We integrate the driver with the mdtest [mdtest] benchmarking tool for HDFS metadata operations.
 Server implementation.
-Each server is implemented in C++ with 3 K LoC and hosts an HDFS (v3.2) namenode (hdf, [n.
-d.]a).
-We use RBF (hdf, [n.
-d.]b) with the HASH_ALL policy, which uses consistent hashing to distribute files evenly across all namenodes for load balancing and creates directories on all namenodes.
-Each server manages a subset of the metadata namespace, connects to its local namenode via the C++ HDFS client library libhdfs3 (lib, [n.
-d.]), and serves client requests while maintaining a path-token map for local hash resolution (§6).
-8. Evaluation
-8.1. Methodology
+Each server is implemented in C++ with 3 K LoC and hosts an HDFS (v3.2) namenode [hdfs].
+We use RBF [hdfsrbf] with the HASH_ALL policy, which uses consistent hashing to distribute files evenly across all namenodes for load balancing and creates directories on all namenodes.
+Each server manages a subset of the metadata namespace, connects to its local namenode via the C++ HDFS client library libhdfs3 [libhdfs3], and serves client requests while maintaining a path-token map for local hash resolution (§VI).
+VIII Evaluation
+VIII-A Methodology
+Table I: Real-world workloads and their metadata operation ratios.
 Workload Operation ratio Read ratio
-Alibaba (Lv et al., 2022) 52.6% open/close, 9.59% create, 3.9% readdir, 69.1%
+Alibaba [lv22] 52.6% open/close, 9.59% create, 3.9% readdir, 69.1%
 0.1% chmod, 11.9% delete, 12.4% stat, 0.2% statdir,
 0.005% mkdir, 0.005% rmdir, 9.3% file rename
-Training (Xu et al., 2024) 54.32% open/close, 28.5% stat, 0.13% readdir, 83.1%
+Training [xu24] 54.32% open/close, 28.5% stat, 0.13% readdir, 83.1%
 9.01% create, 0.13% mkdir, 0.13% rmdir,
 9.01% delete, 0.13% statdir
-Thumb (Xu et al., 2024) 57.01% open/close, 28.44% stat, 0.13% readdir, 85.7%
+Thumb [xu24] 57.01% open/close, 28.44% stat, 0.13% readdir, 85.7%
 14.16% create, 0.13% mkdir, 0.13% statdir
-LinkedIn (Ren et al., 2014) 84% open/getattr, 9% create/mkdir, 84%
+LinkedIn [ren14] 84% open/getattr, 9% create/mkdir, 84%
 7% chmod/delete/rename
-Table 1. Real-world workloads and their metadata operation ratios.
 Testbed.
-We evaluate FMCache on a testbed comprising a 3.2 Tbps two-pipeline Tofino switch (tof, [n.
-d.]; swi, [n.
-d.]) and three physical machines.
+We evaluate Fletch on a testbed comprising a 3.2 Tbps two-pipeline Tofino switch [tofino, switchthpt] and three physical machines.
 Two machines host servers, each with a 2.40 GHz 10-core Intel Xeon Silver 4210R CPU, 128 GiB DRAM, and a 2 TB HDD (Dell PERC H330 Mini).
@@
 The client machine uses one pipeline, and the two server machines use another pipeline.
-All counter arrays (§7) reside in the server-connected pipeline.
-Since our two-pipeline Tofino switch lacks native support for cross-pipeline recirculation, we physically connect the designated ingress pipeline hosting the lock counter arrays to another ingress pipeline using a physical wire (Sheng et al., 2023).
-This enables requests from another ingress pipeline to be recirculated to the designated ingress pipeline (§5).
+All counter arrays (§VII) reside in the server-connected pipeline.
+Since our two-pipeline Tofino switch lacks native support for cross-pipeline recirculation, we physically connect the designated ingress pipeline hosting the lock counter arrays to another ingress pipeline using a physical wire [sheng25].
+This enables requests from another ingress pipeline to be recirculated to the designated ingress pipeline (§V).
 Future programmable switches with native cross-pipeline recirculation would eliminate this requirement.
 Workloads.
-We evaluate FMCache using four real-world workload traces: (i) Alibaba’s Pangu file-system instances (Alibaba) (Lv et al., 2022), (ii) convolutional neural network training (Training) (Xu et al., 2024), (iii) the processing of one million thumbnail images (Thumb) (Xu et al., 2024), and (iv) a LinkedIn HDFS cluster (LinkedIn) (Ren et al., 2014).
-Table 1 summarizes the proportion of metadata operations and the read ratio for each workload.
+We evaluate Fletch using four real-world workload traces: (i) Alibaba’s Pangu file-system instances (Alibaba) [lv22], (ii) convolutional neural network training (Training) [xu24], (iii) the processing of one million thumbnail images (Thumb) [xu24], and (iv) a LinkedIn HDFS cluster (LinkedIn) [ren14].
+Table I summarizes the proportion of metadata operations and the read ratio for each workload.
 We refine the workloads for our evaluation.
-To focus on metadata performance, we exclude reads and writes of file data, following prior studies (Niazi et al., 2017; Li et al., 2017a; Ren et al., 2014; Lv et al., 2022).
-For Training and Thumb that include file reads and writes (Xu et al., 2024), we exclude these operations and normalize the remaining metadata operations.
+To focus on metadata performance, we exclude reads and writes of file data, following prior studies [niazi17, li17, ren14, lv22].
+For Training and Thumb that include file reads and writes [xu24], we exclude these operations and normalize the remaining metadata operations.
 Since we exclude file writes, we treat close as a read operation, while it updates both modification and access timestamps if the closed file has been updated.
-HDFS updates access timestamps hourly by default (def, [n.
-d.]), and we exclude timestamp updates as they are infrequent.
-For LinkedIn, as the original paper (Ren et al., 2014) does not provide operation ratios, we adjust the ratios based on trace analysis (Lv et al., 2022) as follows: open (42%), getattr (42%), create (4.5%), mkdir (4.5%), chmod (1%), delete (3%), and rename (3%).
-We assign the smallest ratio to chmod as it is less frequent than delete and rename (Lv et al., 2022).
+HDFS updates access timestamps hourly by default [defaultsetting], and we exclude timestamp updates as they are infrequent.
+For LinkedIn, as the original paper [ren14] does not provide operation ratios, we adjust the ratios based on trace analysis [lv22] as follows: open (42%), getattr (42%), create (4.5%), mkdir (4.5%), chmod (1%), delete (3%), and rename (3%).
+We assign the smallest ratio to chmod as it is less frequent than delete and rename [lv22].
 Further, getattr corresponds to stat and statdir in HDFS.
-We replace getattr with stat, as file operations are much more frequent than directory operations (Lv et al., 2022), and focus on file renaming for rename.
+We replace getattr with stat, as file operations are much more frequent than directory operations [lv22], and focus on file renaming for rename.
 Generation of metadata operations.
-We use mdtest (mdt, [n.
-d.]) to generate file-system namespaces and metadata operations.
-By default, we configure a path depth of nine, as metadata requests are often aggregated at small depths (Agrawal et al., 2007; Meyer and Bolosky, 2012; Douceur and Bolosky, 1999)5.1).
-We create 32 million empty files to focus on metadata performance; empty files are also used in prior evaluation (Ren et al., 2014; Xiao et al., 2015; Niazi et al., 2017; Li et al., 2017a; Lv et al., 2022; Wang et al., 2023).
-To simulate workload skewness, we use the 80/20 rule, with 80% of operations on 20% of directories and files (Wang et al., 2021a; Xu et al., 2024).
-We model file access frequencies across levels using a power-law distribution with an exponent of 0.9 and randomly assigning an operation type using the ratios in Table 1.
+We use mdtest [mdtest] to generate file-system namespaces and metadata operations.
+By default, we configure a path depth of nine, as metadata requests are often aggregated at small depths [agrawal07, meyer12, douceur99]V-A).
+We create 32 million empty files to focus on metadata performance; empty files are also used in prior evaluation [ren14, xiao15, niazi17, li17, lv22, wang23].
+To simulate workload skewness, we use the 80/20 rule, with 80% of operations on 20% of directories and files [wang21, xu24].
+We model file access frequencies across levels using a power-law distribution with an exponent of 0.9 and randomly assigning an operation type using the ratios in Table I.
 For statdir and readdir, we choose the parent directory of the selected file.
 For mkdir and rmdir, we use separate directories to avoid removing non-empty directories.
-We address various workload settings in §8.3.
+We address various workload settings in §VIII-C.
 We mix all metadata operations in each workload.
-However, rename, delete, and rmdir inherently involve metadata modifications that necessitate the granting and revoking of leases in HDFS (Wang et al., 2009).
+However, rename, delete, and rmdir inherently involve metadata modifications that necessitate the granting and revoking of leases in HDFS [wang09].
 The lease-based operations on frequently created and deleted files can slow down all metadata operations when we perform stress tests.
@@
 Scaling servers.
-Although our testbed contains only two physical servers, our Tofino switch has significantly higher forwarding throughput (3.2 Tbps (swi, [n.
-d.])) than server throughput (tens of KOPS) and is not our evaluation bottleneck.
-To simulate larger-scale deployment with limited hardware while maintaining realism in evaluation, we adopt the well-known server rotation approach as in prior in-switch caching studies (Jin et al., 2017; Sheng et al., 2023).
+Although our testbed contains only two physical servers, our Tofino switch has significantly higher forwarding throughput (3.2 Tbps [switchthpt]) than server throughput (tens of KOPS) and is not our evaluation bottleneck.
+To simulate larger-scale deployment with limited hardware while maintaining realism in evaluation, we adopt the well-known server rotation approach as in prior in-switch caching studies [jin17, sheng25].
 Let $N$ be the number of simulated servers.
-We assign files to $N$ simulated servers using the HDFS’s RBF policy (§7).
+We assign files to $N$ simulated servers using the HDFS’s RBF policy (§VII).
 We calculate each server’s load from file access frequencies, and identify the bottlenecked server with the highest load.
@@
 Baselines.
-We consider client-side caching (CCache) by faithfully following the state-of-the-art client-side caching implementations in IndexFS (Ren et al., 2014) and InfiniFS (Lv et al., 2022)9).
+We consider client-side caching (CCache) by faithfully following the state-of-the-art client-side caching implementations in IndexFS [ren14] and InfiniFS [lv22]IX).
 CCache uses RocksDB (v6.22.1) as a key-value store to keep all metadata instead of in an HDFS namenode in each simulated server to eliminate HDFS’s path resolution overhead.
 Each CCache client locally caches directories’ permission metadata, and forwards all read requests to servers for attribute retrieval.
-We implement lazy invalidation (Lv et al., 2022) for cache consistency, which outperforms lease-based cache management (Ren et al., 2014).
-We do not compare FMCache with in-switch key-value caches (e.g., NetCache (Jin et al., 2017) and FarReach (Sheng et al., 2023)) since they are designed for key-value stores and do not support file-system semantics (i.e., they cannot ensure correctness and consistency for file-system operations).
-We evaluate four schemes: (i) NoCache, which does not employ any caching and performs metadata operations directly with HDFS namenodes in all simulated servers; (ii) CCache, our client-side caching implementation; (iii) FMCache, which extends NoCache with in-switch caching; and (iv) FMCache+, which extends CCache with in-switch caching.
-Before each experiment, we pre-load 32 million files into HDFS namenodes for NoCache and FMCache, and the corresponding paths and metadata into RocksDB for CCache and FMCache+.
-We pre-load the 5,000 hottest files and their ancestors into the in-switch cache for FMCache and FMCache+.
-We set the pre-defined CMS threshold of FMCache and FMCache+ to 10 and 20, respectively (a larger threshold is used for FMCache+ due to its higher throughput with client-side caching), and reset the CMS and the frequency counter array every two seconds.
+We implement lazy invalidation [lv22] for cache consistency, which outperforms lease-based cache management [ren14].
+We do not compare Fletch with in-switch key-value cache systems (e.g., NetCache [jin17] and FarReach [sheng25]) since they are designed for key-value stores and do not support file-system semantics (i.e., they cannot ensure correctness and consistency for file-system operations).
+We evaluate four schemes: (i) NoCache, which does not employ any caching and performs metadata operations directly with HDFS namenodes in all simulated servers; (ii) CCache, our client-side caching implementation; (iii) Fletch, which extends NoCache with in-switch caching; and (iv) Fletch+, which extends CCache with in-switch caching.
+Before each experiment, we pre-load 32 million files into HDFS namenodes for NoCache and Fletch, and the corresponding paths and metadata into RocksDB for CCache and Fletch+.
+We pre-load the 5,000 hottest files and their ancestors into the in-switch cache for Fletch and Fletch+.
+We set the pre-defined CMS threshold of Fletch and Fletch+ to 10 and 20, respectively (a larger threshold is used for Fletch+ due to its higher throughput with client-side caching), and reset the CMS and the frequency counter array every two seconds.
 We simulate 128 client threads, which sufficiently saturate back-end servers.
-The client-side cache of each simulated client in CCache and FMCache+ is allocated 4 MiB (Lv et al., 2022).
+The client-side cache of each simulated client in CCache and Fletch+ is allocated 4 MiB [lv22].
 We plot the average results over five runs, with error bars as 95% confidence intervals under the Student’s t-distribution.
-8.2. Performance Analysis
+VIII-B Performance Analysis
 Refer to caption Refer to caption
@@
 (b) 128 servers
-Figure 7. (Exp#1) Performance under real-world workloads.
+Figure 7: (Exp#1) Performance under real-world workloads.
 (Exp#1) Throughput under real-world workloads.
 Figure 7 shows the throughput results on 16 and 128 simulated servers via server rotation under four real-world workloads.
-Under 16 servers, FMCache increases NoCache’s throughput by 22.9%, 51.7%, 53.7%, and 47.6% in Alibaba, Training, Thumb, and LinkedIn, respectively, due to load balancing via in-switch caching, while NoCache suffers from load imbalance.
-FMCache has lower throughput than CCache by around 38.1% for all workloads.
+Under 16 servers, Fletch increases NoCache’s throughput by 22.9%, 51.7%, 53.7%, and 47.6% in Alibaba, Training, Thumb, and LinkedIn, respectively, due to load balancing via in-switch caching, while NoCache suffers from load imbalance.
+Fletch has lower throughput than CCache by around 38.1% for all workloads.
 The reason is that CCache reduces the path resolution overhead of HDFS by caching directories’ permission metadata in clients and bypasses the overhead of lease management and distributed transactions in HDFS.
-However, by integrating FMCache with CCache, FMCache+ increases CCache’s throughput by 8.0%, 33.5%, 36.3%, and 31.5% in Alibaba, Training, Thumb, and Linkedin, respectively, as CCache forwards all read requests to servers for attribute retrieval (§8.1) while FMCache+ improves load balancing via in-switch caching.
-Our results show that FMCache complements client-side caching to improve metadata performance.
-Under 128 servers, FMCache increases NoCache’s throughput by 11.0%, 134.6%, 181.6%, and 71.2% in Alibaba, Training, Thumb, and LinkedIn, respectively.
-FMCache+ increases CCache’s throughput by 14.7%, 103.6%, 139.6%, and 57.3% for the same workloads, respectively.
-FMCache and FMCache+ achieve significantly higher throughput gains than that in 16 servers (except in Alibaba) as they improve server scalability via load balancing.
+However, by integrating Fletch with CCache, Fletch+ increases CCache’s throughput by 8.0%, 33.5%, 36.3%, and 31.5% in Alibaba, Training, Thumb, and Linkedin, respectively, as CCache forwards all read requests to servers for attribute retrieval (§VIII-A) while Fletch+ improves load balancing via in-switch caching.
+Our results show that Fletch complements client-side caching to improve metadata performance.
+Under 128 servers, Fletch increases NoCache’s throughput by 11.0%, 134.6%, 181.6%, and 71.2% in Alibaba, Training, Thumb, and LinkedIn, respectively.
+Fletch+ increases CCache’s throughput by 14.7%, 103.6%, 139.6%, and 57.3% for the same workloads, respectively.
+Fletch and Fletch+ achieve significantly higher throughput gains than that in 16 servers (except in Alibaba) as they improve server scalability via load balancing.
 The throughput gains from 16 to 128 servers are marginal in Alibaba, which has the largest write ratio among the four workloads.
-Since FMCache and FMCache+ adopt write-through caching, maintaining cache consistency for extensive writes incurs substantial overhead.
+Since Fletch and Fletch+ adopt write-through caching, maintaining cache consistency for extensive writes incurs substantial overhead.
 (Exp#2) Single-operation performance.
 Figure 8 shows the throughput of individual metadata operations: open, stat, create, mkdir, rename, chmod, delete, and rmdir.
-For read operations (open and stat), FMCache increases NoCache’s throughput by 80.1% and 80.5%, while FMCache+ increases CCache’s throughput by 74.9% and 77.5%, respectively.
-For write operations (create, mkdir, rename, chmod, delete, and rmdir), FMCache has lower throughput than NoCache by 2.7%, 0.2%, 13.2%, 36.5%, 12.7%, and 14.6%, respectively, while FMCache+ has lower throughput than CCache, by 4.9%, 3.5%, 7.2%, 12.3%, 8.1%, and 6.4%, respectively.
+For read operations (open and stat), Fletch increases NoCache’s throughput by 80.1% and 80.5%, while Fletch+ increases CCache’s throughput by 74.9% and 77.5%, respectively.
+For write operations (create, mkdir, rename, chmod, delete, and rmdir), Fletch has lower throughput than NoCache by 2.7%, 0.2%, 13.2%, 36.5%, 12.7%, and 14.6%, respectively, while Fletch+ has lower throughput than CCache, by 4.9%, 3.5%, 7.2%, 12.3%, 8.1%, and 6.4%, respectively.
 The performance drops stem from the switch’s cache maintenance overhead.
@@
 In contrast, rename, delete, and rmdir on cached paths only mark the paths as deleted in the in-switch cache, while create and mkdir only involve creations on new files and directories on HDFS namenodes, respectively, and will not trigger cache updates.
-Figure 8. (Exp#2) Single-operation performance.
+Figure 8: (Exp#2) Single-operation performance.
 (Exp#3) Impact of chmod ratio.
-We further analyze chmod, which triggers frequent cache updates and shows the most performance drops in FMCache and FMCache+ (Exp#2).
+We further analyze chmod, which triggers frequent cache updates and shows the most performance drops in Fletch and Fletch+ (Exp#2).
 We generate mixed read-write workloads composed of open (i.e., reads) and chmod (i.e., writes) with different ratios.
 Each of open and chmod follows a power-law distribution with an exponent of 0.9.
-Figure 9 shows that at 0% chmod ratio, FMCache and FMCache+ achieve higher throughput than NoCache and CCache, respectively, but as the chmod ratio increases, their throughput decreases, while the throughput of NoCache and CCache remains stable.
-FMCache and FMCache+ begin to show throughput degradations when the chmod ratio exceeds 50%.
-At 100% chmod ratio, FMCache and FMCache+ reach the lowest throughput, 36.5% and 12.3% lower than NoCache and CCache, respectively.
-Nevertheless, real-world workloads have low chmod ratios (Table 1), so FMCache and FMCache+ still maintain performance gains in practice (Exp#1).
-FMCache’s multi-level read-write locking design (§5) significantly reduces the overhead in writes.
+Figure 9 shows that at 0% chmod ratio, Fletch and Fletch+ achieve higher throughput than NoCache and CCache, respectively, but as the chmod ratio increases, their throughput decreases, while the throughput of NoCache and CCache remains stable.
+Fletch and Fletch+ begin to show throughput degradations when the chmod ratio exceeds 50%.
+At 100% chmod ratio, Fletch and Fletch+ reach the lowest throughput, 36.5% and 12.3% lower than NoCache and CCache, respectively.
+Nevertheless, real-world workloads have low chmod ratios (Table I), so Fletch and Fletch+ still maintain performance gains in practice (Exp#1).
+Fletch’s multi-level read-write locking design (§V) significantly reduces the overhead in writes.
 We compare multi-level locking (MultiLock) with single-level locking (SingleLock), which always maps a full path to the first lock counter array.
-We measure the recirculation count (i.e., number of recirculations per request) by the monitoring tool Barefoot Shell (bfs, [n.
-d.]).
+We measure the recirculation count (i.e., number of recirculations per request) by the monitoring tool Barefoot Shell [bfshell].
 We do not compare MultiLock and locking-disabled in-switch caches, which cannot maintain cache consistency.
-Table 2 shows that for read-only and write-only workloads, the recirculation counts for both SingleLock and MultiLock are 7.63 and 1, respectively, as FMCache recirculates a read request multiple times for path resolution and a write request once for lock access.
+Table II shows that for read-only and write-only workloads, the recirculation counts for both SingleLock and MultiLock are 7.63 and 1, respectively, as Fletch recirculates a read request multiple times for path resolution and a write request once for lock access.
 MultiLock significantly reduces the recirculation count of SingleLock (e.g., by 93.1% for 75% chmod) due to mitigated lock contention.
-Figure 9. (Exp#3) Throughput versus chmod ratio.
+Figure 9: (Exp#3) Throughput versus chmod ratio.
+Table II: (Exp#3) Recirculation count versus chmod ratio.
 chmod ratio SingleLock MultiLock
@@
 100% 1 1
-Table 2. (Exp#3) Recirculation count versus chmod ratio.
 (Exp#4) Latency analysis.
-We analyze request latencies by adjusting the request sending rate to a target throughput, as in prior studies (Cheng et al., 2015; Jin et al., 2017; Sheng et al., 2023; Didona and Zwaenepoel, 2019), so as to analyze the trade-off between latencies and throughput.
-We focus on (i) a read-only workload that issues 32 million open requests and (ii) the Alibaba workload, which has the largest write ratio (Table 1).
-The two workloads show FMCache’s best- and worst-case performance, respectively, under write-through caching (§3.1).
+We analyze request latencies by adjusting the request sending rate to a target throughput, as in prior studies [cheng15, jin17, sheng25, didona19], so as to analyze the trade-off between latencies and throughput.
+We focus on (i) a read-only workload that issues 32 million open requests and (ii) the Alibaba workload, which has the largest write ratio (Table I).
+The two workloads show Fletch’s best- and worst-case performance, respectively, under write-through caching (§III-A).
 We follow a power-law distribution with an exponent of 0.9 and consider 16 simulated servers.
@@
 (f) Alibaba
-Figure 10. (Exp#4) Latency analysis.
+Figure 10: (Exp#4) Latency analysis.
 Figure 10 shows the average, p95, and p99 latency results.
@@
 When the target throughput increases to 0.2 MOPS, NoCache starts to show increasing latencies due to load imbalance under higher loads.
-FMCache maintains low latencies via in-switch caching.
-For example, at target throughout 0.26 MOPS, FMCache reduces NoCache’s average, p95, and p99 latencies by 64.6%, 89.8%, and 87.7%, respectively.
-The same trend is also observed for FMCache+ and CCache.
-For example, at target throughput 0.59 MOPS, FMCache+ reduces CCache’s average, p95, and p99 latencies by 26.9%, 14.7%, and 77.0%, respectively.
-For the Alibaba workload, FMCache also outperforms NoCache.
-For example, at target throughput 0.15 MOPS, FMCache reduces NoCache’s average, p95 and p99 latencies by 66.1%, 65.4%, and 37.2%, respectively.
-For FMCache+ and CCache, at low throughput, they have comparable average and p95 latencies, while FMCache has slightly higher p99 latencies due to the switch’s cache maintenance overhead.
-At high throughput (e.g., 0.35 MOPS), FMCache+ reduces CCache’s average, p95, and p99 latencies by 25.5%, 63.7%, and 15.5%, respectively.
-8.3. Impact of Workload Settings
-We examine FMCache’s robustness across different workload settings.
-In the interest of space, we focus on Alibaba (with the largest write ratio) and Thumb (with the largest read ratio), while Training and LinkedIn show consistent behaviors with Thumb due to their similar read ratios.
+Fletch maintains low latencies via in-switch caching.
+For example, at target throughout 0.26 MOPS, Fletch reduces NoCache’s average, p95, and p99 latencies by 64.6%, 89.8%, and 87.7%, respectively.
+The same trend is also observed for Fletch+ and CCache.
+For example, at target throughput 0.59 MOPS, Fletch+ reduces CCache’s average, p95, and p99 latencies by 26.9%, 14.7%, and 77.0%, respectively.
+For the Alibaba workload, Fletch also outperforms NoCache.
+For example, at target throughput 0.15 MOPS, Fletch reduces NoCache’s average, p95 and p99 latencies by 66.1%, 65.4%, and 37.2%, respectively.
+For Fletch+ and CCache, at low throughput, they have comparable average and p95 latencies, while Fletch has slightly higher p99 latencies due to the switch’s cache maintenance overhead.
+At high throughput (e.g., 0.35 MOPS), Fletch+ reduces CCache’s average, p95, and p99 latencies by 25.5%, 63.7%, and 15.5%, respectively.
+VIII-C Impact of Workload Settings
 (Exp#5) Impact of file access frequency assignment.
@@
 We consider three sorted file sequence: (i) high-level-first (HLF), which sorts files in descending order of their levels (i.e., files at higher levels have higher access frequencies), (ii) low-level-first (LLF), which sorts files in ascending order of their levels (i.e., files at lower levels have higher access frequencies), and (iii) random (our default), which generates a random sequence of files across all levels.
-Figure 11 shows that FMCache and FMCache+ still outperform NoCache and CCache, respectively, under different file access frequency assignments.
-For example, in Thumb (the most read-intensive), FMCache and FMCache+ increase the throughput of NoCache and CCache by 53.7-69.0% and 30.2-44.7%, respectively.
+Figure 11 shows that Fletch and Fletch+ still outperform NoCache and CCache, respectively, under different file access frequency assignments.
+For example, in Thumb (the most read-intensive), Fletch and Fletch+ increase the throughput of NoCache and CCache by 53.7-69.0% and 30.2-44.7%, respectively.
 (Exp#6) Impact of access skewness.
@@
 A larger exponent implies a more skewed access pattern.
-Figure 12 shows that under the uniform workload, FMCache and FMCache+ have slightly less throughput than NoCache and CCache, respectively, by up to 5.0%, since most requests come from uncached paths and are served by servers, while FMCache and FMCache+ incur extra cache maintenance overhead.
+Figure 12 shows that under the uniform workload, Fletch and Fletch+ have slightly less throughput than NoCache and CCache, respectively, by up to 5.0%, since most requests come from uncached paths and are served by servers, while Fletch and Fletch+ incur extra overhead for maintaining cache consistency.
 Refer to caption Refer to caption
 (a) Alibaba
-(b) Thumb
-Figure 11. (Exp#5) Impact of file access frequency assignment.
+(b) Training
+Refer to caption Refer to caption
+(c) Thumb
+(d) LinkedIn
+Figure 11: (Exp#5) Impact of file access frequency assignment.
 Refer to caption Refer to caption
 (a) Alibaba
-(b) Thumb
-Figure 12. (Exp#6) Impact of access skewness.
+(b) Training
+Refer to caption Refer to caption
+(c) Thumb
+(d) LinkedIn
+Figure 12: (Exp#6) Impact of access skewness.
+[…]
+For Alibaba and LinkedIn, as the exponent increases from 0.9 to 1.0, the throughput of Fletch and Fletch+ decreases.
+Alibaba has the largest write ratio, while LinkedIn has the highest chmod ratio.
+They incur significant maintenance overhead for cache consistency.
+Nevertheless, at exponent 1.0 for LinkedIn, Fletch and Fletch+ still increase the throughput of NoCache and CCache by 84.3% and 45.1%, respectively.
+(Exp#7) Impact of maximum path depth.
+We vary the maximum path depth as 3, 5, 7, and 9 (our default).
+Figure 13 shows that Fletch and Fletch+ always outperform NoCache and CCache, respectively.
+For example, in Thumb, Fletch and Fletch+ increase the throughput of NoCache and CCache by 53.7-77.9% and 34.1-36.3%, respectively.
 Refer to caption Refer to caption
 (a) Alibaba
-(b) Thumb
-Figure 13. (Exp#7) Impact of maximum path depth.
-[…]
-For Alibaba, as the exponent increases from 0.9 to 1.0, the throughput of FMCache and FMCache+ decreases.
-The reason is that Alibaba has the largest write ratio and incurs significant cache maintenance overhead.
-Nevertheless, at exponent 1.0, FMCache and FMCache+ still increase the throughput of NoCache and CCache by 37.6% and 25.5%, respectively.
-(Exp#7) Impact of maximum path depth.
-We vary the maximum path depth as 3, 5, 7, and 9 (our default).
-Figure 13 shows that FMCache and FMCache+ always outperform NoCache and CCache, respectively.
-For example, in Thumb, FMCache and FMCache+ increase the throughput of NoCache and CCache by 53.7-77.9% and 34.1-36.3%, respectively.
+(b) Training
+Refer to caption Refer to caption
+(c) Thumb
+(d) LinkedIn
+Figure 13: (Exp#7) Impact of maximum path depth.
 (Exp#8) Impact of dynamic workloads.
 We consider dynamic workloads with varying access frequencies of files over time.
-We follow the prior studies (Li et al., 2016; Jin et al., 2017; Sheng et al., 2023) to generate the hot-in dynamic pattern, which periodically selects the 100 least-frequently accessed files, re-assigns them with the highest access frequencies, and adjusts the access frequencies of other files accordingly to maintain a power-law distribution.
+We follow the prior studies [li16, jin17, sheng25] to generate the hot-in dynamic pattern, which periodically selects the 100 least-frequently accessed files, re-assigns them with the highest access frequencies, and adjusts the access frequencies of other files accordingly to maintain a power-law distribution.
 We set the change period as 20 seconds and run each scheme for 200 seconds to measure per-second throughput.
 We disable server rotation as the system states change under dynamic workloads; instead, we issue workloads to the two physical servers and measure performance directly.
-Figure 14 shows that for Thumb, FMCache and FMCache+ show performance dips due to periodic changes of file access frequencies.
-Before new hot records are admitted, performance dips occur, but FMCache and FMCache+ quickly admit new hot records and return to high performance with path-aware cache management (§4).
-Also, local hash collision resolution (§6) incurs minimal overhead to cache admission and eviction.
-For Alibaba, which has the largest write ratio, FMCache and FMCache+ have marginal performance gains over NoCache and CCache, respectively.
-Nevertheless, FMCache and FMCache+ still effectively respond to dynamic workloads.
+Figure 14 shows that for Training, Thumb, and LinkedIn, Fletch and Fletch+ show performance dips due to periodic changes of file access frequencies.
+Before new hot records are admitted, performance dips occur, but Fletch and Fletch+ quickly admit new hot records and return to high performance with path-aware cache management (§IV).
+Also, local hash collision resolution (§VI) incurs minimal overhead to cache admission and eviction.
+For Alibaba, which has the largest write ratio, Fletch and Fletch+ have marginal performance gains over NoCache and CCache, respectively, as they incur cache consistency overhead.
+Nevertheless, Fletch and Fletch+ still effectively respond to dynamic workloads.
 Refer to caption Refer to caption
 (a) Alibaba
-(b) Thumb
-Figure 14. (Exp#8) Impact of dynamic workloads.
-8.4. Switch Deployment
+(b) Training
+Refer to caption Refer to caption
+(c) Thumb
+(d) LinkedIn
+Figure 14: (Exp#8) Impact of dynamic workloads.
+VIII-D Switch Deployment
 (Exp#9) Switch resource usage.
 We measure the resource usage on the Tofino switch for (i) SRAM (15 MiB), (ii) number of stages (12), (iii) number of ALUs (48), and (iv) PHV size (768 bytes).
-We also quote the resource usage of two in-switch key-value caches, NetCache (Jin et al., 2017) and FarReach (Sheng et al., 2023).
-Table 3 shows that NoCache and CCache use the least resources for L2/L3 forwarding, while FMCache and FMCache+ also support L2/L3 forwarding to process packets unrelated to metadata operations at line rate (3.2 Tbps for the Tofino switch (swi, [n.
-d.])).
-Additionally, FMCache and FMCache+ consume additional resources for caching (e.g., using SRAM and ALUs to cache metadata, track access frequencies, and maintain lock counters, and using PHVs to parse metadata requests), yet their resource usage is comparable to NetCache and FarReach (state-of-the-art in-switch key-value caches).
+We also quote the resource usage of two in-switch key-value cache systems, NetCache [jin17] and FarReach [sheng25], based on the egress switch resource usage reported in the paper [sheng25].
+Table III shows that NoCache and CCache use the least resources for L2/L3 forwarding, while Fletch and Fletch+ also support L2/L3 forwarding to process packets unrelated to metadata operations at line rate (3.2 Tbps for the Tofino switch [switchthpt]).
+Additionally, Fletch and Fletch+ consume additional resources for caching (e.g., using SRAM and ALUs to cache metadata, track access frequencies, and maintain lock counters, and using PHVs to parse metadata requests), yet their resource usage is comparable to NetCache and FarReach (state-of-the-art in-switch key-value cache approaches).
+Table III: (Exp#9) Switch resource usage (numbers in brackets refer to the fractions of used resources over total available ones). Numbers of NetCache and FarReach are quoted from [sheng25].
 Scheme SRAM (KiB) # Stages # ALUs PHV size (bytes)
@@
 CCache 288 (1.9%) 4 (33.3%) 0 (0%) 256 (33.3%)
-NetCache (Jin et al., 2017) 8800 (57.3%) 12 (100%) 45 (93.8%) 528 (68.8%)
-FarReach (Sheng et al., 2023) 8992 (58.5%) 12 (100%) 45 (93.8%) 499 (65.0%)
-FMCache 8976 (58.4%) 12 (100%) 47 (97.6%) 712 (92.7%)
-FMCache+ 8976 (58.4%) 12 (100%) 47 (97.6%) 712 (92.7%)
-Table 3. (Exp#9) Switch resource usage (numbers in brackets refer to the fractions of used resources over total available ones). Numbers of NetCache and FarReach are quoted from (Sheng et al., 2023).
-9. Related Work
+NetCache [jin17] 7856 (51.2%) 12 (100%) 45 (93.8%) 528 (68.8%)
+FarReach [sheng25] 8080 (52.6%) 12 (100%) 45 (93.8%) 499 (65.0%)
+Fletch 8976 (58.4%) 12 (100%) 47 (97.6%) 712 (92.7%)
+Fletch+ 8976 (58.4%) 12 (100%) 47 (97.6%) 712 (92.7%)
+IX Related Work
 Scaling file-system metadata management.
 Metadata partitioning distributes file-system management across multiple servers for scalability.
-There are two primary approaches: (i) dynamic sub-tree partitioning, which distributes namespace sub-trees across servers (Weil et al., 2006; Sevilla et al., 2015; Wang et al., 2021a), and (ii) hash-based partitioning, which distributes file-system metadata across servers via hashing (Patil and Gibson, 2011; Thomson and Abadi, 2015; Niazi et al., 2017; Pan et al., 2021; Liao and Abadi, 2023; Wang et al., 2023).
-FMCache complements these approaches as an in-switch cache that absorbs operations upstream of file-system metadata layers.
+There are two primary approaches: (i) dynamic sub-tree partitioning, which distributes namespace sub-trees across servers [weil06, sevilla15, wang21], and (ii) hash-based partitioning, which distributes file-system metadata across servers via hashing [patil11, thomson15, niazi17, pan21, liao23, wang23].
+Fletch complements these approaches as an in-switch cache that absorbs operations upstream of file-system metadata layers.
 Client-side caching.
-PanFS (Welch et al., 2008) caches file and directory metadata and provides callbacks for cache consistency.
-IndexFS (Ren et al., 2014) and LocoFS (Li et al., 2017a) use lease-based caching for metadata management and invalidate cache entries upon lease expiration, but incur high overhead for renewing cache entries’ leases.
-InfiniFS (Lv et al., 2022) applies lazy invalidation for directory access metadata to limit the overhead of lease-based caching.
+PanFS [welch08] caches file and directory metadata and provides callbacks for cache consistency.
+IndexFS [ren14] and LocoFS [li17] use lease-based caching for metadata management and invalidate cache entries upon lease expiration, but incur high overhead for renewing cache entries’ leases.
+InfiniFS [lv22] applies lazy invalidation for directory access metadata to limit the overhead of lease-based caching.
 Client-side caching often incurs high client-side complexity and overhead in maintaining cache consistency across a large number of clients.
-FMCache simplifies cache consistency by caching file-system metadata in a programmable switch that lies on the critical paths of multiple clients.
+Fletch simplifies cache consistency by caching file-system metadata in a programmable switch that lies on the critical paths of multiple clients.
 In-switch caching.
-Programmable switches have been extensively studied for concurrency control (Li et al., 2017b; Zhu et al., 2019), network monitoring (Gupta et al., 2018), replication coordination (Jin et al., 2018), remote procedure calls (Zhao et al., 2023), key-value stream aggregation (He et al., 2023), and distributed lock management (Yu et al., 2020; Zhang et al., 2024).
+Programmable switches have been extensively studied for concurrency control [li17eris, zhu19], network monitoring [gupta18], replication coordination [jin18], remote procedure calls [zhao23], key-value stream aggregation [he23], and distributed lock management [yu20, zhang24].
 Several studies explore in-switch caching.
-SwitchKV (Li et al., 2016) and NetCache (Jin et al., 2017) design write-through in-switch caching for read-intensive workloads, and DistCache (Liu et al., 2019) designs distributed write-through caching across multiple switches.
-Pegasus (Li et al., 2020) and TurboKV (Eldakiky et al., 2021) cache replica-to-server mappings for replica selection.
-Mind (Lee et al., 2021) caches object-to-memory mappings for disaggregated memory systems.
-Concordia (Wang et al., 2021b) tracks cache copy locations and states to address concurrency in shared memory systems.
-FarReach (Sheng et al., 2023) designs fault-tolerant write-back caching for write-intensive workloads.
+SwitchKV [li16] and NetCache [jin17] design write-through in-switch caching for read-intensive workloads, and DistCache [liu19] designs distributed write-through caching across multiple switches.
+Pegasus [li20] and TurboKV [eldakiky20] cache replica-to-server mappings for replica selection.
+Mind [lee21] caches object-to-memory mappings for disaggregated memory systems.
+Concordia [wang21concordia] tracks cache copy locations and states to address concurrency in shared memory systems.
+FarReach [sheng25] designs fault-tolerant write-back caching for write-intensive workloads, and its extension, DistReach [sheng25], supports multiple switches.
 The above in-switch caches are designed for key-value stores, which differ from file-system semantics.
-AsyncFS (Xu et al., 2024) proposes in-switch tracking of directory updates, while maintaining a client-side metadata cache for path resolution.
-In contrast, FMCache moves file-system metadata caching to switches, and also complements client-side caching.
-10. Conclusion
-FMCache is an in-switch file-system metadata caching framework, aiming to achieve high throughput and load balancing for distributed file-system metadata services.
+SwitchFS [xu24] proposes in-switch tracking of directory updates, while maintaining a client-side metadata cache for path resolution.
+In contrast, Fletch moves file-system metadata caching to switches, and also complements client-side caching.
+X Conclusion
+Fletch is an in-switch file-system metadata caching framework, aiming to achieve high throughput and load balancing for distributed file-system metadata services.
 It employs path-aware cache management, multi-level read-write locking, and local hash collision resolution.
-Experiments on a Tofino-switch testbed show that FMCache achieves significant throughput gains and complements client-side caching.
+Experiments on a Tofino-switch testbed show that Fletch achieves significant throughput gains and complements client-side caching.